### v1

# Gromov’s Compactness Theorem for the Intrinsic Timed Hausdorff Distance

Abstract.
We prove Gromov’s Compactness Theorem for the Intrinsic timed Hausdorff ($\tau-H$) convergence of timed-metric-spaces using timed-Fréchet maps. Our proof introduces the notion of ”addresses” and provides a new way of stating Gromov’s original compactness theorem for Gromov-Hausdorff (GH) convergence of metric spaces. We also obtain a new Arzela-Ascoli Theorem for real valued uniformly bounded Lipschitz functions on GH converging compact metric spaces. The intrinsic timed-Hausdorff distance between timed-metric-spaces was first defined by Sakovich-Sormani to define a weak notion of convergence for space-times, and our compactness theorem will soon be applied to advance their work in this direction.
R. Perales was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/EFP6]. M. Che was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32]. The research was also funded in part by Sormani’s PSC-CUNY and NSF DMS-1612409 grants.
1. Introduction
One of the most profound theorems in metric geometry is Gromov’s Compactness Theorem: a sequence of equibounded and equicompact metric spaces has a subsequence converging in the Gromov-Hausdorff sense to a compact metric space. See Gromov’s presentation in [16] and his application of a stronger version to groups of polynomial growth in [15]. We review these notions and state this theorem precisely in Sections 2.12.3.
When a sequence of metric spaces converges in the Gromov-Hausdorff sense then one can control distances on the limit spaces. Not only can one control distances on these converging sequences of metric spaces, but one can also prove Arzela-Ascoli theorems for sequences of uniformly bounded Lipchitz functions on the sequence of spaces (c.f. Greene-Petersen [14]).
Before Gromov’s Compactness Theorem, Riemannian Geometers could only study the limits of sequences of diffeomorphic Riemannian manifolds. With Gromov-Hausdorff convergence, geometers can study sequences of Riemannian manifolds with different topologies and even different dimensions. A sequence of compact Riemannian manifolds can be canonically converted into compact metric spaces using the Riemannian distance. With uniform upper diameter bounds and lower Ricci curvature bounds, Gromov proved that Riemannian manifolds are equicompact and equibounded and thus any sequence subconverges to a limit space [16][17]. This was applied with great success by Cheeger, Colding, Fukaya, Naber, Sormani, Wei and others in many papers including [10][13][8][9][30][11].
Lorentzian geometers would also like to study sequences of smooth Lorentzian manifolds which are not diffeomorphic. In [25], Sakovich and Sormani introduced the notion of the intrinsic timed-Hausdorff convergence as one possible way to achieve this goal. They canonically converted a sequence of smooth causally-null-compactifiable space-times into timed-metric-spaces, $(X_{j},d_{j},\tau_{j})$, which are metric spaces, $(X_{j},d_{j})$, endowed with bounded Lipschitz one time functions, $\tau_{j}:X_{j}\to[0,\tau_{max}]$. They then defined the intrinsic timed-Hausdorff convergence of these timed-metric-spaces. We review these definitions in Sections 2.52.7.
In this paper we prove a compactness theorem for this notion of timed-Hausdorff convergence: a sequence of equibounded and equicompact timed-metric-spaces with uniformly bounded time functions converges in the intrinsic timed-Hausdorff sense. For the full somewhat stronger statement, see Theorem 3.1.
Our compactness theorem is proven in Sections 3.13.3. In Section 3.1 we begin the proof by imitating Gromov’s proof of his compactness theorem in [15]. As we added significant details to his original proof, we developed the notion of “addresses” for points in Section 3.2. With these “addresses” we complete a stronger statement of the compactness theorem which controls the convergence of time in a uniform way in Section 3.3.
Our notion of “addresses” provides a new way of stating Gromov’s Compactness Theorem for sequences of metric spaces, that we state and prove in Section 4.1. We also restate an Arzela-Ascoli Theorem for sequences of real valued uniformly bounded Lipschitz functions on a converging sequence of metric spaces using “addresses” in Section 4.2. We suggest a more general Arzela-Ascoli Conjecture that could be proven using “addresses” in Section 4.3.
Sakovich-Sormani introduced the notion of the intrinsic timed-Hausdorff distance in [25] to define a new notion of convergence for smooth Lorentzian space-times. They describe a method of converting smooth causally null compactifiable space-times into compact timed metric spaces in a canonical way using the cosmological time function of Anderson-Galloway-Howard [4] and the null distance of Sormani-Vega [29]. Although we do not wish to go into the details of this conversion in this paper, we point out that our compactness theorem is immediately a precompactness theorem for sequences of such smooth Lorentzian manifolds producing limit spaces that are timed-metric-spaces in [25]. Further applications of our compactness theorem to Lorentzian manifolds will appear in future work.
It should be noted that some people have used the Sormani-Vega null distance of [29] to convert space-times into metric spaces, without keeping track of the time functions. See the work of Allen, Burtscher, García-Heveling, Kunzinger, and Sämann in [2, 3, 6, 19]. If one does this, then one can apply Gromov’s original compactness theorem to obtain a timeless GH limit space which is only a metric space. However, one needs a time function to define the causal structure on the limit space [25]. We believe the work of Allen, Burtscher, and García-Heveling should be extendable to the stronger notion of timed-Hausdorff convergence studied here.
There are special classes of space-times where the cosmological time functions are easy to control. These include the big-bang space-times and future developed space-times. Sormani-Vega [28] and Sakovich-Sormani [25] defined convergence for such space-times based on pointed GH convergence of Gromov [17] and GH Convergence for Metric Pairs introduced by Che, Galaz-García, Guijarro, and Membrillo Solis in [7]. Such notions already have compactness theorems proven by Gromov [17] and by Che and Ahumada Gómez [1] respectively. Sakovich-Sormani defined the intrinsic timed-Hausdorff distance to apply to a far wider class of space-times including sub-space-times that exhaust asymptotically flat space-times [25] and we need our new compactness theorem to apply in these settings.
Other notions of weak convergence of space-times do not involve the convergence of the space-times into metric spaces or timed-metric-spaces. Minguzzi-Suhr [21] have a compactness result in the setting of “Lorentzian metric spaces” (which are not metric spaces) building upon work of Noldus [23, 24]. See also the work of Mondino–Sämann [22] in the setting of “Lorentzian pre-length spaces”, as introduced in [18], where they define a notion of Lorentzian Gromov–Hausdorff convergence and obtain a precompactness principle using causal diamonds and Lorentz distances. The spaces considered in these works are not metric spaces. It is not yet clear how their notions of convergence relate to the Sakovich-Sormani notions defined using timed-metric-spaces in [25]. This is worth exploring further.
Finally we note that, in upcoming work, Sakovich-Sormani will be introducing timed intrinsic flat convergence of space-times whose limit spaces are rectifiable timed-metric-spaces with biLipchitz charts [26]. Our compactness theorem proven in this paper will be applied in that paper in an essential way.
Acknowledgements: We would like to thank Anna Sakovich (Uppsala) for online conversations in July 2025 that lead us to pursuing this project. We all gratefully acknowledge support from the Simons Center for Geometry and Physics, Stony Brook University where we had the opportunity to meet together in person in September 2025 as part of the Program Geometry and Convergence in Mathematical General Relativity. R. Perales was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/EFP6]. M. Che was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32]. C. Sormani was funded in part by a PSC-CUNY grant. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
2. Background
2.1. Review of the Hausdorff Distance
Recall that the Hausdorff distance between sets, $U,W\subset Z$, is extrinsic in that it depends on the distance in the extrinsic space:
The Hausdorff distance between subsets, $U,W\subset Z$, lying in a common metric space, $(Z,d_{Z})$, id defined
(1) $d_{H}^{Z}\bigg(U,W\bigg)=\inf\left\{r>0:\begin{array}[]{c}\forall u\in U,\exists w\in W\textrm{ s.t. }d_{Z}(u,w)<r\\ \forall w\in W,\exists u\in U\textrm{ s.t. }d_{Z}(u,w)<r\end{array}\right\}.$
2.2. Review of Gromov-Hausdorff Distance
Gromov defined his intrinsic Hausdorff (GH) distance between distinct metric spaces which do not lie in a common metric space in [17]. This is now called the Gromov-Hausdorff distance. The GH distance is intrinsic in the sense that it depends only on the intrinsic geometry of the two given metric spaces, $(X_{j},d_{j})$, and not the extrinsic geometry of some common ambient space.
In order to define this distance intrinsically, Gromov considered all possible common metric spaces, $(Z,d_{Z})$, and applied maps: $\varphi_{j}:X_{j}\to Z$ that are distance preserving
(2) $d_{Z}(\varphi_{j}(p),\varphi_{j}(q))=d_{j}(p,q)\qquad\forall p,q\in X_{j}.$
and takes the infimum:
(3) $d_{GH}\bigg((X_{1},d_{1}),(X_{2},d_{2})\bigg)=\inf d_{H}^{Z}\bigg(\varphi_{1}(X_{1}),\varphi_{2}(X_{2})\bigg).$
Thanks to the infimum taken over all possible distance preserving maps, the Gromov-Hausdorff distance depends only on the intrinsic geometry of the original pair of metric spaces not on any particular external metric space, $(Z,d_{Z})$. Yet at the same time the GH distance between metric spaces has many of the same properties as the Hausdorff distance. In particular there are correspondences between $X_{1}$ and $X_{2}$ with controlled distance distortion that are often used to define the GH distance, but we will not use correspondences in this paper.
2.3. Review of Gromov’s Compactness Theorem
Recall Gromov’s Compactness Theorem: proven in [17] and [15]:
Theorem 2.1 (Gromov Compactness Theorem).
If $(X_{j},d_{j},\tau_{j})$ are a sequence of compact metric spaces that are equibounded,
(4) $\exists D>0\,s.t.\,\operatorname{diam}_{d_{j}}(X_{j})\leq D,$
equicompact,
(5) $\forall R\in(0,D]\,\,\exists N(R)\in\mathbb{N}\,s.t.\,\exists x^{j}_{R,i}\in X_{j}\,\,s.t.\,\,X_{j}\subset\bigcup_{i=1}^{N(R)}B(x^{j}_{R,i},R),$
then a subsequence converges in the Gromov-Hausdorff sense to a compact metric space $(\bar{X}_{\infty},d_{\infty})$:
(6) $d_{GH}((X_{j},d_{j}),(X_{\infty},d_{\infty}))\to 0.$
In fact, there exist a uniform compact metric space, $(Z,d_{Z})$, and distance preserving maps,
(7) $\varphi_{j}:X_{j}\to Z,$
such that the Hausdorff distance between the images converges to zero:
(8) $d_{H}^{Z}\bigg(\varphi_{j}(X_{j}),\varphi_{\infty}(X_{\infty})\bigg)\to 0.$
ADD MORE???
2.4. Review of Fréchet Maps
Here we review the notion of a Fréchet map [12].  [^1] In [25] this was accidentally attributed to Kuratowski [20] and will be fixed.
First recall that a separable metric-space, $(X,d)$, is a metric-space containing a countable dense collection of points, $\{x_{1},x_{2},...\}\subset X$.
Recall also the Banach space
[…]
where
(10) $d_{\ell^{\infty}}((s_{1},s_{2},...),(r_{1},r_{2},...))=\sup\{|s_{i}-r_{i}|\,:\,i\in{\mathbb{N}}\}.$
Definition 2.2.
Given a separable metric-space, $(X,d)$, for any countable dense collection of points $\mathcal{N}=\{x_{1},x_{2},$...\} of $X$, one can define a Fréchet map
(11) $\kappa_{X}=\kappa_{X,\mathcal{N}}:(X,d)\to(\ell^{\infty},d_{\ell^{\infty}})$
by
[…]
 [^2] In v1 of [25] this was accidentally attributed to Kuratowski [20] and will be fixed.
The following theorem has been credited to Fréchet:
Theorem 2.3 (Fréchet).
Fréchet maps as in Definition 2.2 are distance preserving:
(13) $d_{\ell^{\infty}}(\kappa_{X}(x),\kappa_{X}(y))=d(x,y)\quad\forall\,x,y\in X.$
It is well known that one can define a version of the Gromov-Hausdorff distance by taking the infimum of the Hausdorff distance between images of Fréchet maps into $\ell^{\infty}$.
Definition 2.4.
The Fréchet-GH distance
(14) $d_{\kappa-GH}((X,d_{X}),(Y,d_{Y}))=\inf d_{H}^{\ell^{\infty}}(\kappa_{X}(X),\kappa_{Y}(Y))$
where the infimum is over all pairs of Fréchet maps
[…]
which is an infimum over all selections of countably dense points and reorderings of these selected points in $X$ and in $Y$.
By Fréchet’s Theorem, one immediately has
(16) $d_{\kappa-GH}((X,d_{X}),(Y,d_{Y}))\geq d_{GH}((X,d_{X}),(Y,d_{Y})).$
In fact this notion is biLipschitz equivalent to the Gromov-Hausdorff distance:
(17) $d_{GH}((X,d_{X}),(Y,d_{Y}))\leq d_{\kappa-GH}((X,d_{X}),(Y,d_{Y}))\leq 2d_{GH}((X,d_{X}),(Y,d_{Y}))$
for any pair of compact metric-spaces. This close relationship with the Gromov-Hausdorff distance is discussed and reproven in the background of [25] and directly inspired their definition of the intrinsic timed Hausdorff distance using timed-Fréchet maps.
2.5. Review of Timed-Metric-Spaces
The notion of a timed-metric-space was introduced by Sakovich-Sormani in [25] because they had constructed a canonical way to convert certain classes of smooth Lorentzian space-times into timed-metric-spaces keeping track of the causal structure of the original Lorentzian manifold. For our purposes, we need only the abstract notion of a timed-metric-space:
Definition 2.5.
A timed-metric-space, $(X,d,\tau)$, is a metric space, $(X,d)$, with a definite, symmetric map $d:X\times X\to[0,\infty),$ satisfying the triangle inequality, that is endowed with a Lipschitz one time function, $\tau:X\to[0,\infty).$ It has a causal structure defined as follows:
(18) $p\in J^{+}(q)\Longleftrightarrow\tau(p)-\tau(q)=d(p,q).$
Note that a map $F:(X,d_{X},\tau_{X})\to(Y,d_{Y},\tau_{Y})$ is distance preserving iff
(19) $d_{Y}(F(p),F(q))=d_{X}(p,q)\qquad\forall p,q\in X.$
It is time preserving iff
(20) $\tau_{Y}(F(p))=\tau_{X}(p)\qquad\forall p\in X.$
It preserves the causal structure iff
(21) $F(p)\in J^{+}_{Y}(F(q))\Longleftrightarrow p\in J^{+}_{X}(q)$
which happens when $F$ is both distance and time preserving.
In this paper we focus on the geometry of timed-metric-spaces and the intrinsic timed Hausdorff distance. Our work applies to study Lorentzian manifolds and their intrinsic timed Hausdorff limits, but we will not be exploring that relationship within this paper. We will not refer further to smooth manifolds.
2.6. Review of Timed-Fréchet Maps
Sakovich-Sormani introduced the following notion in [25, Definition 4.21]:
Definition 2.6.
Given a compact timed-metric-space, $(X,d,\tau)$, and a countably dense collection of points,
(22) ${\mathcal{N}}=\{x_{1},x_{2},x_{3},...\}\subset X,$
we define a timed-Fréchet map $\kappa_{\tau,X}=\kappa_{\tau,X,{\mathcal{N}}}:X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty}$, as follows:
[…]
 [^3] In v1 of [25] these were accidentally named timed-Fréchet maps and this will be fixed.
For the reader’s convenience, we state here propositions 4.24 and 4.25 from [25].
Proposition 2.7.
Any timed-Fréchet map, $\kappa_{\tau,X}:X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty},$ is distance preserving,
(24) $d_{\ell^{\infty}}(\kappa_{\tau,X}(x),\kappa_{\tau,X}(y))=d(x,y),$
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau:X\to[0,\tau_{max}]$ and any choice of countably dense points in $X$.
Proposition 2.8.
If we consider the infinite dimensional timed-metric-space
(25) $([0,\tau_{max}]\times\ell^{\infty},d_{\ell^{\infty}},\tau_{\infty}),$
where
(26) $\tau_{\infty}(w_{0},w_{1},w_{2},...)=w_{0},$
endowed with a causal structure defined by
(27) $w\in J^{+}(z)\Longleftrightarrow\tau_{\infty}(w)-\tau_{\infty}(z)=d_{\ell^{\infty}}(w,z),$
then every timed-Fréchet map $\kappa_{\tau,X}:X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty}$ is time preserving:
(28) $\tau_{\infty}(\kappa_{\tau,X}(x))=\tau(x)$
and preserves causality:
(29) $\kappa_{\tau,X}(p)\in J^{+}(\kappa_{\tau,X}(q))\Longleftrightarrow p\in J^{+}(q)$
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau:X\to[0,\tau_{max}]$ that encodes causality as in (18).
2.7. Review of the Intrinsic Timed Hausdorff Distance
In [25, Definition 4.22] Sakovich and Sormani introduced the following notion within their definition of the intrinsic timed Hausdorff distance between causally null compactifiable space-times:
Definition 2.9.
The intrinsic timed Hausdorff distance between two compact timed-metric-spaces, $(X_{i},d_{i},\tau_{i})$, $i=1,2$, is given by
(30) $d^{\kappa}_{\tau-H}\Big((X_{1},d_{1},\tau_{1}),(X_{2},d_{2},\tau_{2})\Big)=\inf\,d^{\ell^{\infty}}_{H}(\kappa_{\tau_{1},X_{1}}(X_{1}),\kappa_{\tau_{2},X_{2}}(X_{2}))$
where the infimum is taken over all possible timed-Fréchet maps, $\kappa_{\tau_{1},X_{1}}:X_{1}\to\ell^{\infty}$ and $\kappa_{\tau_{2},X_{2}}:X_{2}\to\ell^{\infty}$ which are found by considering all countable dense collections of points in the $X_{i}$ and all reorderings of these collections of points.
They proved this notion of distance is definite in [25, Theorem 5.14], which we state below.
Theorem 2.10.
The intrinsic timed Hausdorff distance between two compact timed-metric-spaces, as in definition 2.9, is definite in the sense that:
(31) $d^{\kappa}_{\tau-H}\Big((X,d_{X},\tau_{X}),(Y,d_{Y},\tau_{Y})\Big)=\inf\,d^{\ell^{\infty}}_{H}(\kappa_{\tau_{X},X}(X),\kappa_{\tau_{Y},Y}(Y))=0$
if and only if there is a distance and time preserving bijection
(32) $F:(X,d_{X},\tau_{X})\to(Y,d_{Y},\tau_{Y}).$
3. Proving the Compactness Theorem with Addresses
In this section we prove our compactness theorem:
Theorem 3.1 (Timed Gromov Compactness Theorem).
If $(X_{j},d_{j},\tau_{j})$ are a sequence of compact timed metric spaces that are equibounded,
(33) $\exists D>0\,s.t.\,\operatorname{diam}_{d_{j}}(X_{j})\leq D,$
equicompact,
(34) $\forall R\in(0,D]\,\exists N(R)\in\mathbb{N}\,s.t.\,\exists\{x^{j}_{R,i}:i=1,\ldots,N(R)\}\,\,s.t.\,\,X_{j}\subset\bigcup_{i=1}^{N(R)}B(x^{j}_{R,i},R)$
and have uniform bounds on their Lipschitz one functions
(35) $\tau_{j}(X_{j})\subset[0,\tau_{max}],$
then a subsequence converges in the intrinsic timed Hausdorff sense to a compact timed metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. In fact, there exist distance and time preserving timed-Fréchet maps,
(36) $\varphi_{j}:X_{j}\to[0,\tau_{max}]\times Z\subset\ell^{\infty},$
where $Z$ is compact, such that
(37) $d_{H}^{[0,\tau_{max}]\times Z}(\varphi_{j}(X_{j}),\varphi_{\infty}(X_{\infty}))\to 0.$
In addition there is an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(38) ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}\textrm{ and }{\mathcal{I}}^{\infty}:{\mathcal{A}}\to\bar{X}_{\infty}$
such that we have uniform convergence as $j\to\infty$,
(39) $\sup_{\alpha\in{\mathcal{A}}}\|\varphi_{j}({\mathcal{I}}^{j}(\alpha))-\varphi_{\infty}({\mathcal{I}}^{\infty}(\alpha))\|_{\infty}\to 0.$
In particular, we have uniform convergence of time,
(40) $\sup_{\alpha\in{\mathcal{A}}}|\tau_{j}({\mathcal{I}}^{j}(\alpha))-\tau_{\infty}({\mathcal{I}}^{\infty}(\alpha))|\to 0,$
and uniform convergence of distances,
(41) $\sup_{\alpha,\alpha^{\prime}\in{\mathcal{A}}}|d_{j}({\mathcal{I}}^{j}(\alpha),{\mathcal{I}}^{j}(\alpha^{\prime}))-d_{\infty}({\mathcal{I}}^{\infty}(\alpha),{\mathcal{I}}^{\infty}(\alpha^{\prime}))|\to 0.$
3.1. Selection of Balls in a Compact Metric Space
Recall the definitions of equiboundedness and equicompactness in the statement of Theorem 2.1. The following proposition is from the first part of Gromov’s proof of his Compactness Criterion in [15] that we will apply later to prove our Theorem 3.1.
In the first proposition, the idea is to carefully select a countable dense collection of points covering each metric space and indexing them in a way which keeps track of where each point is located. Part of the statement is taken exactly from Gromov but we add in the precise indexing of the points $x_{a}$ and provide more details in the proof. See Figure 3.1. We will apply this indexing later to define “addresses”.
Proposition 3.2.
Suppose $X_{j}$ is a metric space satisfying the equiboundedness and equicompactness conditions as in the statement of Theorem 2.1. Take the sequence $\varepsilon_{i}=2^{-i}$ and let $N_{i}=N(\varepsilon_{i})$ be natural numbers such that $X_{j}$ can be covered by $N_{i}$ balls of radius $\varepsilon_{i}$.
Then we can find a countable dense subset
(42) $\{x_{a}^{j}:\,a\in A\}\subset X_{j}$
using a uniform index set $A$, not depending on $j$, as described below. For each $i\in{\mathbb{N}}$, we let $A_{i}$ be the finite set,
(43) $A_{i}=\Bigl\{(a_{1},a_{2},\ldots,a_{i}):\,1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2},\dots,1\leq a_{i}\leq N_{i}\Bigr\},$
and we define the projection maps,
(44) $p_{i}:A_{i+1}\to A_{i}\qquad p_{i}:(a_{1},\ldots,a_{i+1})\mapsto(a_{1},\ldots,a_{i}).$
Consider the countable set
(45) $A=\bigsqcup_{i=1}^{\infty}A_{i}.$
to be the index set for our countable collection of points with the maps
(46) $I^{j}_{i}:A_{i}\to X_{j}\quad\textrm{ and }\quad I^{j}:A\to X_{j},$
such that
(47) $I^{j}(a)=I^{j}_{i}(a)\in X_{j}\qquad\forall a\in A_{i}\subset A,$
and such that the points,
(48) $x^{j}_{a}=I^{j}(a)=I^{j}_{i}(a)\qquad\forall a\in A_{i},$
are centers of well chosen balls of radius $\varepsilon_{i}$ in $X_{j}$ for $a\in A_{i}$.
The collection of points
(49) ${\mathcal{N}}_{j}=I^{j}(A)=\{x_{a}^{j}:\,a\in A\}=\bigcup_{i\in{\mathbb{N}}}I^{j}_{i}(A_{i})\subset X_{j}$
is countably dense in $X_{j}$ and has been carefully indexed to satisfy the following two properties. First, $I^{j}_{i}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, which means
(50) $X_{j}\subset\bigcup_{a\in A_{i}}B_{d_{j}}(x^{j}_{a},\varepsilon_{i})$
Second, for each $a\in A_{i+1}$,
(51) $I^{j}_{i+1}(a)\in B_{d_{j}}(I^{j}_{i}(p_{i}(a)),2\varepsilon_{i})\subset X_{j}$
which means
(52) $x^{j}_{(a_{1},...a_{i},a_{i+1})}\in B_{d_{j}}(x^{j}_{(a_{1},...a_{i})},2\varepsilon_{i})$
for any $a=(a_{1},\dots,a_{i+1})\in A_{i+1}$.
Figure 3.1. An index set with $N(1)=1,N(2)=3,N(3)=4$ and the corresponding points in $X_{j}$.
Proof.
We will complete the proof of the construction and of (50) and (52) by induction on $i\in{\mathbb{N}}$. Note that, once this induction is completed, we will see that the set $I^{j}(A)$ is dense in $X_{j}$ because we can take any $x\in X_{j}$ and any $\delta>0$, and choose $i$ sufficiently large that $\varepsilon_{i}<\delta$, and then apply (50) to see that,
(53) $\exists x^{j}_{a}\in I^{j}(A)\,\,s.t.\,\,d_{j}(x,x^{j}_{a})<\varepsilon_{i}<\delta.$
By construction, the set $A$ and its image $I^{j}(A)$ are countable because they are countable unions of finite sets, $A_{i}$, and their images, $I^{j}_{i}(A_{i})$, respectively. Thus we need only complete the inductive proof to prove this proposition.
We begin the induction argument with $i=1$. By equiboundedness and the definition of $N_{1}=N(\varepsilon_{1})$, we have an $\varepsilon_{1}$-net of points,
(54) $\{x^{j}_{(1)},x^{j}_{(2)},x^{j}_{(3)},...x^{j}_{(N_{1})}\}\subset X$
indexed by
(55) $A_{1}=\{(1),(2),\ldots,(N_{1})\}$
such that
(56) $X_{j}\subset\bigcup_{a\in A_{1}}B_{d_{j}}(x^{j}_{a},\varepsilon_{1})=\bigcup_{a_{1}=1}^{N_{1}}B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})$
which gives us (50) for $i=1$.
By equiboundedness and the definition of $N_{2}=N(\varepsilon_{2})$, for each ball of radius $\varepsilon_{1}$ (indexed by $a_{1}=1,2,...,N_{1}$) as above, we can choose $N_{2}$ points in $X_{j}$ such that
(57) $B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})\subset\bigcup_{a_{2}=1}^{N_{2}}B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2})$
with
(58) $B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})\cap B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2})\neq\emptyset\qquad a_{2}=1,\ldots,N_{2}.$
Thus
(59) $x^{j}_{(a_{1},a_{2})}\in B_{d_{j}}(x^{j}_{(a_{1})},2\varepsilon_{1})\subset X_{j},$
So we have an index set,
(60) $A_{2}=\{(a_{1},a_{2}):1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2}\},$
and a projection map
(61) $p_{1}:A_{2}\to A_{1}\qquad p_{1}:(a_{1},a_{2})\mapsto(a_{1})$
and the map $I^{j}_{2}:A_{2}\to X_{j}$ identifying the center points,
(62) $x^{j}_{(a_{1},a_{2})}=I^{j}_{2}((a_{1},a_{2}))$
of the balls of radius $\varepsilon_{2}$. Observe that by (59) and the definition of the projection map, $p_{1}$, we have indexed the points
(63) $x^{j}_{(a_{1},a_{2})}\in B_{d_{j}}(x^{j}_{p_{1}(a_{1},a_{2})},2\varepsilon_{1})\subset X_{j},$
which gives (52) for $i=1$. This completes the base step of our proof by induction.
For further understanding, note that by combining (56) and (57) we have
$X_{j}\subset\bigcup_{a_{1}=1}^{N_{1}}B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})\subset\bigcup_{a_{1}=1}^{N_{1}}\bigcup_{a_{2}=1}^{N_{2}}B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2})=\bigcup_{a\in A_{2}}B_{d_{j}}(x^{j}_{a},\varepsilon_{2})$
which gives us (50) for $i=2$.
Now we proceed with the induction step. Assume we already have a collection of points indexed by the sets
$A_{1}$ $=$ $\{(a_{1}):1\leq a_{1}\leq N_{1}\},$
$A_{2}$ $=$ $\{(a_{1},a_{2}):1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2}\},$
$A_{3}$ $=$ $\{(a_{1},a_{2},a_{3}):1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2},\,1\leq a_{3}\leq N_{3}\},$
$\vdots$
$A_{i}$ $=$ $\{(a_{1},a_{2},\ldots,a_{i}):\,1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2},\ldots,1\leq a_{i}\leq N_{i}\},$
with projection maps,
(64) $p_{1}\colon A_{2}\to A_{1},\quad p_{2}\colon A_{3}\to A_{2},\quad\ldots,\quad p_{i-1}\colon A_{i}\to A_{i-1},$
and maps
(65) $I^{j}_{1}\colon A_{1}\to X_{j},\quad I^{j}_{2}\colon A_{2}\to X_{j},\quad\ldots,\quad I^{j}_{i}\colon A_{i}\to X_{j},$
identifying the center points,
$x^{j}_{(a_{1},a_{2})}\,\,=$ $I^{j}_{2}((a_{1},a_{2}))$ $\in\,\,B_{d_{j}}(x^{j}_{(a_{1})},2\varepsilon_{1}),$
$x^{j}_{(a_{1},a_{2},a_{3})}\,\,=$ $I^{j}_{3}((a_{1},a_{2},a_{3}))$ $\in\,\,B_{d_{j}}(x^{j}_{(a_{1},a_{2})},2\varepsilon_{2}),$
$\vdots$
$x^{j}_{(a_{1},a_{2},...,a_{i})}\,\,=$ $I^{j}_{i}((a_{1},a_{2},...,a_{i}))$ $\in\,\,B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i-1})},2\varepsilon_{i-1}),$
such that
$B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})$ $\subset$ $\bigcup_{a_{2}=1}^{N_{2}}B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2}),$
$B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2})$ $\subset$ $\bigcup_{a_{3}=1}^{N_{3}}B_{d_{j}}(x^{j}_{(a_{1},a_{2},a_{3})},\varepsilon_{3}),$
$\vdots$
$B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i-1})},\varepsilon_{i-1})$ $\subset$ $\bigcup_{a_{i}=1}^{N_{i}}B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i-1},a_{i})},\varepsilon_{i}).$
By equicontinuity, each ball of radius $\varepsilon_{i}$ centered at
(66) $x^{j}_{(a_{1},a_{2},...,a_{i})}=I^{j}_{i}((a_{1},a_{2},...,a_{i})),$
can be covered by $N_{i+1}$ balls of radius $\varepsilon_{i+1}$ centered at points
(67) $x^{j}_{(a_{1},a_{2},...,a_{i},a_{i+1})}\,\,\in\,\,B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i-1},a_{i})},2\varepsilon_{i}),$
where $a_{i+1}\in\{1,2,...,N_{i+1}\}$. That is
(68) $B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i})},\varepsilon_{i})\,\,\subset\,\,\bigcup_{a_{i+1}=1}^{N_{i+1}}B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i},a_{i+1})},\varepsilon_{i+1}).$
So we have all these new points together indexed by the new index set,
(69) $A_{i+1}=\{(a_{1},a_{2},\ldots,a_{i+1}):1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2},\ldots,1\leq a_{i+1}\leq N_{i+1}\},$
and new maps $I^{j}_{i+1}:A_{i+1}\to X_{j}$ defined by
(70) $x^{j}_{(a_{1},a_{2},...,a_{i},a_{i+1})}=I^{j}_{i+1}((a_{1},a_{2},...,a_{i},a_{i+1})).$
The new projection map,
(71) $p_{i}:A_{i+1}\to A_{i},$
with (67) gives (52) and with (68) gives (50) inductively. ∎
3.2. The Index Set of Addresses
A natural consequence of our description of Gromov’s Proposition 3.2, is that we have a uniform way of providing “addresses” identifying all points in equicompact collections of metric spaces. See Figure 3.2. We make this precise in the following proposition:
Proposition 3.3.
Given $A$ as in Proposition 3.2, we define an address, $\alpha\in{\mathcal{A}}$, to be a sequence,
(72) $\alpha=\{\alpha_{1},\alpha_{2},\alpha_{3},...\}\subset A$
such that
(73) $\alpha_{i}=(a_{1},a_{2},...a_{i})\in A_{i}\subset A\qquad\forall i\in{\mathbb{N}}$
so that
(74) $p_{i}(\alpha_{i+1})=\alpha_{i}\qquad\forall i\in{\mathbb{N}}.$
Let
(75) $q_{i}:{\mathcal{A}}\to A_{i}\textrm{ s.t. }q_{i}(\alpha)=\alpha_{i}.$
Then for any $X_{j}$ and any maps $I^{j}_{i}\colon A_{i}\to X_{j}$ satisfying the conclusions, (46)- (51), of Proposition 3.2, the map, ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}$, defined by
(76) ${\mathcal{I}}^{j}(\alpha)=\lim_{i\to\infty}I^{j}_{i}(q_{i}(\alpha))$
is well defined and surjective. That is, every point $x\in X_{j}$, has an address,
(77) $\alpha=\alpha^{j}(x)=\{\alpha^{j}_{1}(x),\alpha^{j}_{2}(x),...\}\in{\mathcal{A}},$
such that $x={\mathcal{I}}^{j}(\alpha^{j}(x))$. In fact,
(78) $x\in\bar{B}_{d_{j}}(I^{j}_{i}(\alpha^{j}_{i}(x)),\varepsilon_{i})\qquad\forall i\in{\mathbb{N}}.$
We do not claim this address is unique. Moreover, we get the same result even if (50) and (51) hold for closed balls rather than open balls.
Figure 3.2. An address $\alpha$ identifying a sequence of points in $X_{j}$ approaching the point ${\mathcal{I}}^{j}(\alpha)$.
Proof.
To prove ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}$ is well defined, we need only show the sequence of points,
(79) $\{I^{j}_{i}(\alpha_{i}):\,i\in\mathbb{N}\}$
is Cauchy. By $p_{i}(\alpha_{i+1})=\alpha_{i}$, and the properties of $I^{j}_{i}$ proven in Proposition 3.2, we have
(80) $I^{j}_{i+1}(\alpha_{i+1})\in\bar{B}_{d_{j}}(I^{j}_{i}(p_{i}(\alpha_{i+1})),2\varepsilon_{i})=\bar{B}_{d_{j}}(I^{j}_{i}(\alpha_{i}),2\varepsilon_{i}).$
Thus for all $i\in{\mathbb{N}}$, we have
(81) $d_{j}(I^{j}_{i}(\alpha_{i}),I^{j}_{i+1}(\alpha_{i+1}))\leq 2\varepsilon_{i}=1/2^{i-1}.$
Applying the triangle inequality and then this estimate repeatedly we have
$d_{j}(I^{j}_{i}(\alpha_{i}),I^{j}_{i+k}(\alpha_{i+k}))$ $\leq$ $d_{j}(I^{j}_{i}(\alpha_{i}),I^{j}_{i+1}(\alpha_{i+1}))+\cdots+d_{j}(I^{j}_{i+k-1}(\alpha_{i+k-1}),I^{j}_{i+k}(\alpha_{i+k}))$
$\leq$ $\tfrac{1}{2^{i-1}}+\cdots+\tfrac{1}{2^{i+k-2}}<\tfrac{1}{2^{i}}=\varepsilon_{i}.$
So the sequence is Cauchy. Since $X_{j}$ is compact, it is complete and so the limit in (76) exists defining ${\mathcal{I}}^{j}(\alpha)$.
Next we show surjectivity, taking any $x\in X_{j}$. Since $X_{j}$ is covered by balls of radius $\varepsilon_{1}$ we know there exists $(a^{j}_{1}(x))\in A_{1}$ such that
(82) $x\in\bar{B}_{d_{j}}(I^{j}_{1}(a^{j}_{1}(x)),\varepsilon_{1}).$
This ball is covered by balls of radius $\varepsilon_{2}$, with centers indexed by $(a_{1},a_{2})\in A_{2}$ so there exists $a^{j}_{2}(x)\in\{1,\ldots,N_{2}\}$ such that
(83) $x\in\bar{B}_{d_{j}}(I^{j}_{2}((a^{j}_{1}(x),a^{j}_{2}(x))),\varepsilon_{2}).$
Continuing forward we select each $a^{j}_{i}(x)$ such that
(84) $x\in\bar{B}_{d_{j}}(I^{j}_{i}((a^{j}_{1}(x),a^{j}_{2}(x),...,a^{j}_{i}(x))),\varepsilon_{i}).$
Taking
(85) $\alpha^{j}(x)=\{\alpha^{j}_{1}(x),\alpha^{j}_{2}(x),...\}\in\mathcal{A}$
where
(86) $\alpha^{j}_{i}(x)=(a^{j}_{1}(x),a^{j}_{2}(x),...,a^{j}_{i}(x))$
gives us (78). By the definition of ${\mathcal{I}}^{j}$, and then by (78), we have
(87) ${\mathcal{I}}^{j}(\alpha^{j}(x))=\lim_{i\to\infty}I^{j}_{i}(\alpha^{j}_{i}(x))=x.$
3.3. Compactness Theorem with Addresses
We can now prove our compactness theorem:
Proof of Theorem 3.1.
The proof begins as in Gromov’s Compactness Criterion in [15] using the equicompactness as we have stated in Proposition 3.2 to define a countable index set $A$ defined as the disjoint union of finite sets,
(88) $A=A_{1}\sqcup A_{2}\sqcup A_{3}\sqcup$...,
and a countable dense collection of points
(89) $I^{j}(A)=\{x^{j}_{a}=I^{j}(a):\,a\in A\}\subset X_{j}$
such that each $I^{j}(A_{i})$ is a finite $\varepsilon_{i}=1/2^{i}$ net for $X_{j}$ and has all the other properties described in Proposition 3.2.
Then we follow the proof of Burago, Burago and Ivanov [5] of Gromov’s Compactness Theorem (to construct a limit space). By the equiboundedness, for any $a,b\in A$, we define
(90) $d^{j}_{a,b}=d_{j}(x^{j}_{a},x^{j}_{b})\in[0,D].$
This is symmetric
(91) $d^{j}_{a,b}=d^{j}_{b,a}\qquad\forall a,b\in A,$
and satisfies the triangle inequality
(92) $d^{j}_{a,c}\leq d^{j}_{a,b}+d^{j}_{b,c}\qquad\forall a,b,c\in A.$
By the equiboundedness of time, for any $a\in A$, we define
(93) $\tau^{j}_{a}=\tau_{j}(x^{j}_{a})\in[0,\tau_{max}].$
Since the time functions are Lipschitz one, we have
(94) $|\tau^{j}_{a}-\tau^{j}_{b}|\leq d^{j}_{a,b}\qquad\forall a,b\in A.$
By the countability of the index set $A$, we take subsequences of subsequences and diagonalize so that for the final subsequence (which we still denote with $j$), we have for each $a\in A$, convergence of the real numbers:
(95) $\lim_{j\to\infty}\tau^{j}_{a}=\tau^{\infty}_{a}\in[0,\tau_{max}].$
We take further subsequences and diagonalize further so that for the final subsequence (which we still denote with $j$), we have for each $a,b\in A$,
(96) $\lim_{j\to\infty}d^{j}_{a,b}=d^{\infty}_{a,b}\in[0,D].$
Taking the limit of (94) we know
(97) $|\tau^{\infty}_{a}-\tau^{\infty}_{b}|\leq d^{\infty}_{a,b}\qquad\forall a,b\in A.$
We now define a quasi metric space,
(98) $(A,d_{\infty})\textrm{ with }d_{\infty}(a,b)=d^{\infty}_{a,b}\in[0,D]$
which satisfies the triangle inequality and is symmetric but is not necessarily definite. By (97) we know that
(99) $d^{\infty}(a,b)=0\,\,\Longrightarrow\,\,\tau^{\infty}_{a}=\tau^{\infty}_{b}.$
Thus we can mod out by points of zero distance apart to define a timed metric space which is definite:
(100) $(X_{\infty},d_{\infty},\tau_{\infty})=(\{[a]:a\in A\},d_{\infty},\tau_{\infty})$
where the points are equivalence classes,
(101) $a]=\{b\in A:\,d_{\infty}(a,b)=0\},$
the distance function is
(102) $d_{\infty}([a],[b])=d^{\infty}_{a,b}\in[0,D],$
and the time function is
(103) $\tau_{\infty}([a])=\tau^{\infty}_{a}\in[0,\tau_{\max}],$
which is well defined by (99) and is Lipschitz one by (97).
We next take the metric completion
(104) $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$
which has the usual definite distance, $d_{\infty}$, extended from $X_{\infty}$ to $\bar{X}_{\infty}$ satisfying symmetry and the triangle inequality, with
(105) $\operatorname{diam}_{d_{\infty}}(\bar{X}_{\infty})\leq D,$
and a bounded time function, $\tau_{\infty}$, extended as a Lipschitz one function from $X_{\infty}$ to $\bar{X}_{\infty}$.
We claim that $\bar{X}_{\infty}$ is compact. Note that to prove this claim is enough to show that for each $i\in\mathbb{N}$,
(106) $\Bigl\{[a]\,:\,a\in A_{i}\Bigr\}\qquad\text{is an $\varepsilon_{i}$-net of $X_{\infty}$.}$
Take $b]\in X_{\infty}$. Thus, there exist $k\in\mathbb{N}$ such that $b\in A_{k}$. Since for each $j\in\mathbb{N}$ we know that $I_{i}^{j}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, by the pigeon-hole principle, there exists $a\in A_{i}$ such that
(107) $d_{j}(I_{i}^{j}(a),I_{i}^{k}(b))\leq\varepsilon_{i}$
for infinitely many indices $i$. Thus, up to a subsequence,
(108) $d_{\infty}([b],[a])=\lim_{j\to\infty}d_{j}(x_{a}^{j},x_{b}^{j})\leq\varepsilon_{i}.$
Hence, we have (106).
Up to here we have followed the proof of Burago, Burago and Ivanov [5] of Gromov’s Compactness Theorem, with corresponding additions to address the fact that we have time functions, to construct a compact timed-metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. Since we will show timed Hausdorff convergence,
(109) $(X_{j},d_{j},\tau_{j})\stackrel{{\scriptstyle\tau-\textrm{H}}}{{\longrightarrow}}(\bar{X}_{\infty},d_{\infty},\tau_{\infty}),$
using timed-Fréchet maps as in Definition 2.9, from now on, we proceed in a different manner to [5].
Note that we can define timed-Fréchet maps,
(110) $\kappa_{\tau_{j},{\mathcal{N}}_{j}}:X_{j}\to[0,\tau_{max}]\times\prod_{a\in A}[0,D]\subset\ell^{\infty}$
by
(111) $\kappa_{\tau_{j},{\mathcal{N}}_{j}}(x)=(\tau_{j}(x),\kappa_{{\mathcal{N}}_{j}}(x))$
where
(112) $\kappa_{{\mathcal{N}}_{j}}(x)\in\prod_{a\in A}[0,D]$
has components
(113) $(\kappa_{\mathcal{N}_{j}}(x))_{a}=d_{j}(I^{j}(a),x).$
To define a timed-Fréchet map for $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$, recall that we have a countable index set, $A$, which is the countable disjoint union of finite index sets, $A_{i}$. We define
(114) $I^{\infty}_{i}:A_{i}\to X_{\infty}\subset\bar{X}_{\infty}\textrm{ and }I^{\infty}:A\to X_{\infty}\subset\bar{X}_{\infty},$
by
(115) $I^{\infty}(a)=I^{\infty}_{i}(a)=[a]\in X_{\infty}\subset\bar{X}_{\infty}\qquad\forall\,a\in A_{i}\subset A,$
and we denote these points as
(116) $x^{\infty}_{a}=I^{\infty}(a)=I^{\infty}_{i}(a).$
Thus,
(117) ${\mathcal{N}}_{\infty}=I^{\infty}(A)=\{x_{a}^{\infty}:\,a\in A\}$
is a countable ordered collection of points, that is dense in $\bar{X}_{\infty}$, since as a set it equals $X_{\infty}$. From this, and recalling (102) and (103), we also have a timed-Fréchet map $\kappa_{\tau_{\infty},\mathcal{N_{\infty}}}$ satisfying (110)-(113) for $j=\infty$.
Before proving (37) we first show that the uniform estimate (39) holds. As a preliminary step we note that we have uniform convergence on finite unions of the $A_{i}$: Recall that by (90), (96) and (102), for each $a,b\in A$, we have pointwise convergence,
(118) $d_{\infty}(x^{\infty}_{a},x^{\infty}_{b})=\lim_{j\to\infty}d^{j}(x^{j}_{a},x^{j}_{b})\in[0,D].$
By (93), (95) and (103) for each $a\in A$, we have pointwise convergence,
(119) $\tau_{\infty}(x^{\infty}_{a})=\lim_{j\to\infty}\tau_{j}(x^{j}_{a}).$
Now, since each $A_{i}$ is finite, we do obtain uniform convergence on finite unions of the $A_{i}$:
(120) $\forall\epsilon>0,\exists N_{\epsilon}^{i}\in{\mathbb{N}},$
such that for all $j\geq N_{\epsilon}^{i}$ we have
(121) $\max\{|\tau_{\infty}(x^{\infty}_{a})-\tau_{j}(x^{j}_{a})|,|d_{\infty}(x^{\infty}_{a},x^{\infty}_{b})-d^{j}(x^{j}_{a},x^{j}_{b})|\}<\epsilon$
where the maximum is over
(122) $a,b\in A_{1}\cup...\cup A_{i}.$
To get better control than just the pointwise convergence as in (118) and (119), we will use Proposition 3.3 to provide an address, $\alpha$, for every point in each $X_{j}$ and $\bar{X}_{\infty}$.
Note that we have proven that $\bar{X}_{\infty}$ satisfies the conclusions (46)-(49), of Proposition 3.2. It satisfies 50 by (106). Moreover, it easily satisfies (51) as follows. Let $a\in A_{i+1}$,
(123) $I^{\infty}_{i+1}(a)\in B_{d_{\infty}}(I^{\infty}_{i}(p_{i}(a)),2\varepsilon_{i})\subset\bar{X}_{\infty}$
We know that for each $j\in\mathbb{N}$, we know that
(124) $d_{j}(x^{j}_{a},x^{j}_{p_{i}(a)})\leq 2\varepsilon_{i}.$
Hence, in the limit, we have
(125) $d_{\infty}(x^{\infty}_{a},x^{\infty}_{p_{i}(a)})=\lim_{j\to\infty}d_{j}(x^{j}_{a},x^{j}_{p_{i}(a)})\leq 2\varepsilon_{i},$
which is the desired estimate.
Since we have applied Proposition 3.2 for each $X_{j}$, $j\in\mathbb{N}$, and the previous paragraph, for each $j\in\mathbb{N}\sqcup\{\infty\}$ there exist surjective maps, ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}$, $j\in\mathbb{N}\cup\{\infty\}$ defined by
(126) ${\mathcal{I}}^{j}(\alpha)=\lim_{i\to\infty}I^{j}_{i}(q_{i}(\alpha))$
that satisfies the conditions of Proposition 3.3.
We claim that for any $\varepsilon>0$ there exist $N_{\varepsilon}\in\mathbb{N}$ such that for all $j\geq N_{\varepsilon}$ we have for any address $\alpha$,
(127) $|\tau_{j}({\mathcal{I}}^{j}(\alpha))-\tau_{\infty}({\mathcal{I}}^{\infty}(\alpha))|\leq\varepsilon$
and that additionally for any $a\in A$, we also have
(128) $|\kappa_{j,a}({\mathcal{I}}^{j}(\alpha))-\kappa_{\infty,a}({\mathcal{I}}^{\infty}(\alpha))|\leq\varepsilon.$
Let $\epsilon=\varepsilon/3$ and $i\in\mathbb{N}$ such that $\varepsilon_{i}<\varepsilon/3$. Then there exists $N_{\epsilon}^{i}\in{\mathbb{N}}$, as in (120), such that for all $j\geq N_{\epsilon}^{i}$ we have (121) for all $j\geq N_{\varepsilon}^{i}$.
We obtain the first inequality using the triangle inequality,
(129) $|\tau_{j}({\mathcal{I}}^{j}(\alpha))-\tau_{\infty}({\mathcal{I}}^{\infty}(\alpha))|\,\,\leq\,\,$ $|\tau_{j}({\mathcal{I}}^{j}(\alpha))-\tau_{j}(I^{j}_{i}(\alpha_{i}))|$
(130) $+|\tau_{j}(I^{j}_{i}(\alpha_{i}))-\tau_{\infty}(I^{j}_{i}(\alpha_{i}))|$
(131) $+|\tau_{\infty}({\mathcal{I}}^{\infty}(\alpha))-\tau_{\infty}(I^{\infty}_{i}(\alpha_{i}))|$
(132) $\,\,\leq\,\,$ $\,\,d_{j}({\mathcal{I}}^{j}(\alpha),I^{j}_{i}(\alpha_{i}))$
(133) $+|\tau_{j}(I^{j}_{i}(\alpha_{i}))-\tau_{\infty}(I^{j}_{i}(\alpha_{i}))|$
(134) $+d_{\infty}({\mathcal{I}}^{\infty}(\alpha),I^{\infty}_{i}(\alpha_{i}))$
(135) $\leq\,\,$ $\,\,\varepsilon_{i}+\epsilon+\varepsilon_{i}\,\,\leq\,\,\varepsilon\qquad\forall j\geq N^{i}_{\epsilon},\quad\quad\quad$
where for the second and third term in the right hand side we used that $\tau_{j},\tau_{\infty}$ are Lipschitz one and (78), and for the term in the middle we use (121). Similarly, we get the second inequality,
(136) $|\kappa_{j,a}({\mathcal{I}}^{j}(\alpha))-\kappa_{\infty,a}({\mathcal{I}}^{\infty}(\alpha))|\,\,\leq\,\,$ $|\kappa_{j,a}({\mathcal{I}}^{j}(\alpha))-\kappa_{j,a}(I^{j}_{i}(\alpha_{i}))|$
(137) $+\,\,|\kappa_{j,a}(I^{j}_{i}(\alpha_{i}))-\kappa_{\infty,a}(I^{\infty}_{i}(\alpha_{i}))|$
(138) $+\,\,|\kappa_{\infty,a}({\mathcal{I}}^{\infty}(\alpha))-\kappa_{\infty,a}(I^{\infty}_{i}(\alpha_{i}))|$
(139) $\,\,\leq\,\,$ $d_{j}({\mathcal{I}}^{j}(\alpha),I^{j}_{i}(\alpha_{i}))$
(140) $+\,\,|\kappa_{j,a}(I^{j}_{i}(\alpha_{i}))-\kappa_{\infty,a}(I^{\infty}_{i}(\alpha_{i}))|$
(141) $+\,\,d_{\infty}({\mathcal{I}}^{\infty}(\alpha),I^{\infty}_{i}(\alpha_{i}))$
(142) $\,\,\leq\,\,$ $\,\,\varepsilon_{i}+\epsilon+\varepsilon_{i}\leq\varepsilon\qquad\forall j\geq N^{i}_{\epsilon},\quad\quad\quad$
where we used that $\kappa_{j},\kappa_{\infty}$ are Lipschitz one, (78) and (121).
In particular, we have uniform convergence
(143) $\tau_{j}\circ{\mathcal{I}}^{j}\to\tau_{\infty}\circ{\mathcal{I}}^{\infty}\textrm{ on }{\mathcal{A}}$
and
(144) $\kappa_{j,a}\circ{\mathcal{I}}^{j}\to\kappa_{\infty,a}\circ{\mathcal{I}}^{\infty}\textrm{ on }{\mathcal{A}}.$
Hence, (40) and (39) hold, where $\varphi_{j}=(\tau_{j},\kappa_{j})$ and $\varphi_{\infty}=(\tau_{\infty},\kappa_{\infty})$.
Now the correspondences
(145) $\mathcal{C}_{j}=\{(\kappa_{\tau_{j},\mathcal{N}_{j}}({\mathcal{I}}^{j}(\alpha)),\kappa_{\tau_{\infty},\mathcal{N}_{\infty}}({\mathcal{I}}^{\infty}(\alpha))):\alpha\in\mathcal{A}\}$
between $\kappa_{\tau_{j},\mathcal{N}_{j}}(X_{j})$ and $\kappa_{\tau_{\infty},\mathcal{N}_{\infty}}(\bar{X}_{\infty})$, for all $j\in\mathbb{N}$, and the uniform convergence (39) imply that
(146) $d^{\ell^{\infty}}_{H}(\kappa_{\tau_{j},\mathcal{N}_{j}}(X_{j}),\kappa_{\tau_{\infty},\mathcal{N}_{\infty}}(\bar{X}_{\infty}))\leq\|\varphi_{j}({\mathcal{I}}^{j}(\alpha))-\varphi_{\infty}({\mathcal{I}}^{\infty}(\alpha))\|_{\infty}\to 0.$
Finally, let us define
(147) $Z=\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})\cup\bigcup_{j\in\mathbb{N}}\kappa_{\mathcal{N}_{j}}(X_{j})\subset\prod_{a\in A}[0,D],$
so $\kappa_{\tau_{j},\mathcal{N}_{j}}\colon X_{j}\to[0,\tau_{max}]\times Z$ for all $j\in\mathbb{N}\cup\{\infty\}$.
We claim that $Z\subset\ell^{\infty}$ is compact. Let $\{z_{j}\}\subset Z$. If there exists $J\in\mathbb{N}$ such that an infinite number of the $z_{j}$’s is contained in $\kappa_{\mathcal{N}_{J}}(X_{J})$ then by compactness of $\kappa_{\mathcal{N}_{J}}(X_{J})$ there is a further subsequence that converges to a point in $\kappa_{\mathcal{N}_{J}}(X_{J})\subset Z$. Otherwise, assume by passing to a subsequence that $z_{j}\in\kappa_{\mathcal{N}_{j}}(X_{j})$ for all $j\in\mathbb{N}$. By (146), for each $j\in\mathbb{N}$ there exists $w_{j}\in\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})$ such that
(148) $d_{\ell^{\infty}}(z_{j},w_{j})\to 0.$
Since we have shown that $\bar{X}_{\infty}$ is compact, there is a subsequence $\{w_{j_{k}}\}$ that converges to a point $w\in\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})\subset Z$. It follows that $\{z_{j_{k}}\}$ converges to $w$. This completes the proof of $Z$ being compact, and finishes the proof of the theorem. ∎
4. Applications of Addresses
In this section we restate Gromov’s Compactness Theorem and an Arzela-Ascoli Theorem with addresses and state an open question.
4.1. Gromov’s Compactness Theorem with Addresses
Theorem 4.1 (Gromov Compactness Theorem with Addresses).
If $(X_{j},d_{j})$ are a sequence of compact metric spaces that are equibounded as in (4) and equicompact as in (5) then a subsequence converges in the Gromov-Hausdorff sense to a compact metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. In fact, there exist distance preserving Fréchet maps,
(149) $\kappa_{j}:X_{j}\to Z\subset\ell^{\infty},$
where $Z$ is compact, such that
(150) $d_{H}^{Z}(\kappa_{j}(X_{j}),\kappa_{\infty}(X_{\infty}))\to 0.$
In addition there is an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(151) ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}\textrm{ and }{\mathcal{I}}^{\infty}:{\mathcal{A}}\to\bar{X}_{\infty}$
such that we have uniform convergence as $j\to\infty$,
(152) $\sup_{\alpha\in{\mathcal{A}}}\|\kappa_{j}({\mathcal{I}}^{j}(\alpha))-\kappa_{\infty}({\mathcal{I}}^{\infty}(\alpha))\|_{\infty}\to 0.$
In particular,
(153) $\sup_{\alpha,\alpha^{\prime}\in{\mathcal{A}}}|d_{j}({\mathcal{I}}^{j}(\alpha),{\mathcal{I}}^{j}(\alpha^{\prime}))-d_{\infty}({\mathcal{I}}^{\infty}(\alpha),{\mathcal{I}}^{\infty}(\alpha^{\prime}))|\to 0.$
This theorem follows exactly as in the proof of our Theorem 3.1 simply by removing all references to the time function. For completeness of exposition, we prove it as a trivial corollary of our theorem:
Proof.
Take $\tau_{j}$ to be constant zero valued function. Then we satisfy the hypotheses of Theorem 3.1. So the conclusions hold. Finally apply
(154) $\varphi_{j}(x)=(\tau_{j}(x),\kappa_{j}(x))=(0,\kappa_{j}(x)).$
4.2. Arzela-Ascoli Theorem with Addresses
Theorem 4.2 (Arzela-Ascoli-with-Addresses).
Given a sequence of compact metric spaces:
(155) $(X_{j},d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(X_{\infty},d_{\infty})$
with a compact limit space, and given uniformly bounded $K$-Lipschitz functions for some $K>0$,
(156) $F_{j}:X_{j}\to[0,F_{max}],$
then a subsequence, also denoted $F_{j}$, converges to a bounded $K$-Lipschitz function
(157) $F_{\infty}:X_{\infty}\to Y_{\infty}$
in the following sense. There exist distance preserving Fréchet maps,
(158) $\kappa_{j}:X_{j}\to Z\subset\ell^{\infty},$
where $Z$ is compact, such that
(159) $d_{H}^{Z}(\kappa_{j}(X_{j}),\kappa_{\infty}(X_{\infty}))\to 0.$
In addition there is an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(160) ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}\textrm{ and }{\mathcal{I}}^{\infty}:{\mathcal{A}}\to\bar{X}_{\infty}$
such that we have uniform convergence as $j\to\infty$, of the images of the address set under the Fréchet maps,
(161) $\sup_{\alpha\in{\mathcal{A}}}\|\kappa_{j}({\mathcal{I}}^{j}(\alpha))-\kappa_{\infty}({\mathcal{I}}^{\infty}(\alpha))\|_{\infty}\to 0,$
and under the uniformly bounded Lipschitz functions
(162) $\sup_{\alpha\in{\mathcal{A}}}|F_{j}({\mathcal{I}}^{j}(\alpha))-F_{\infty}({\mathcal{I}}^{\infty}(\alpha))|\to 0.$
Proof.
Gromov proved that with a compact GH limit, we have an equibounded and equicompact sequence of metric spaces [16]. Setting
(163) $\tau_{j}(p)=F_{j}(p)/K\qquad\forall p\in X_{j},$
we have timed metric spaces, $(X_{j},d_{j},\tau_{j})$ satisfying the hypotheses of Theorem 3.1. The conclusions immediately follow from that theorem and re-applying (163). ∎
4.3. Arzela-Ascoli Conjecture with Addresses
It would be interesting to apply the address method to more precisely describe the following stronger Arzela-Ascoli Theorem:
Conjecture 4.3.
Given a pair of Gromov-Hausdorff converging sequences of compact metric space:
(164) $X_{j}\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}X_{\infty}\textrm{ and }Y_{j}\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}Y_{\infty}$
and given uniformly bounded $K$-Lipschitz functions for some $K>0$,
(165) $F_{j}:X_{j}\to Y_{j},$
then a subsequence converges to a bounded $K$-Lipschitz function
(166) $F_{\infty}:X_{\infty}\to Y_{\infty}.$
A proof of this theorem can be found in various places including [27]. However a new proof using the addresses would be able to describe the convergence of the functions more precisely as a kind of uniform convergence.