### v2

# Gromov’s Compactness Theorem for the Intrinsic Timed-Hausdorff Distance

Abstract.
The intrinsic timed-Hausdorff distance between timed-metric-spaces, first introduced by Sakovich–Sormani, yields a weak notion of convergence for space-times. In this paper we prove a compactness theorem for the intrinsic timed-Hausdorff convergence of timed-metric-spaces using timed-Fréchet maps. Our proof introduces the notion of “addresses” and provides a new way of stating Gromov’s original compactness theorem for Gromov–Hausdorff convergence of metric spaces. We also obtain a new Arzelà–Ascoli theorem for real valued uniformly bounded Lipschitz functions on Gromov–Hausdorff converging compact metric spaces.
R. Perales was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/EFP6]. M. Che was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32]. The research was also funded in part by Sormani’s PSC-CUNY and NSF DMS-1612409 grants.
1. Introduction
One of the most profound theorems in metric geometry is Gromov’s Compactness Theorem, which states that a sequence of equibounded and equicompact metric spaces has a subsequence converging in the Gromov–Hausdorff sense to a compact metric space [16, 17]. Moreover, when a sequence of metric spaces converges in the Gromov–Hausdorff sense one can prove Arzelà–Ascoli theorems for sequences of uniformly bounded Lipchitz functions on the sequence of spaces (c.f. Greene–Petersen [15]).
A sequence of compact Riemannian manifolds can be canonically converted into compact metric spaces using the Riemannian distance. With Gromov–Hausdorff convergence, one can study sequences of Riemannian manifolds with different topologies and even different dimensions. With uniform upper diameter bounds and lower Ricci curvature bounds, Gromov proved that Riemannian manifolds are equicompact and equibounded and thus any sequence subconverges to a limit space [17, 18]. This was applied with great success by Cheeger, Colding, Fukaya, Naber, Sormani, Wei and others in many papers including [8, 9, 10, 11, 14, 31].
One would also like to study sequences of smooth Lorentzian manifolds which are not diffeomorphic. In [26], Sakovich and Sormani introduced the notion of the intrinsic timed-Hausdorff convergence as one possible way to achieve this. They described a method of converting smooth causally null compactifiable space-times into compact timed-metric-spaces in a canonical way using the cosmological time function of Anderson–Galloway–Howard [4] and the null distance of Sormani–Vega [30]. In general, timed-metric-spaces, $(X_{j},d_{j},\tau_{j})$, are metric spaces, $(X_{j},d_{j})$, endowed with a Lipschitz one  [^1] In this manuscript we refer to Lipschitz functions with Lipschitz constant bounded above by $1$ as Lipschitz one function, $\tau_{j}\colon X_{j}\to\mathbb{R}$, and a causal structure. They then defined the intrinsic timed-Hausdorff convergence of these timed-metric-spaces.
In this paper we prove a compactness theorem for the notion of intrinsic timed-Hausdorff convergence. For the full, somewhat stronger statement, see Theorem 3.1. Note that no knowledge of Lorentzian geometry is required throughout this manuscript, but further applications of our compactness theorem to Lorentzian manifolds will appear in future work.
Theorem 1.1 (Timed Gromov Compactness Theorem–short version).
If $(X_{j},d_{j},\tau_{j})$ is a sequence of compact timed metric spaces that are equibounded,
(1) $\exists D>0\,s.t.\,\operatorname{diam}_{d_{j}}(X_{j})\leq D,$
equicompact,
(2) $\forall R\in(0,D]\,\exists N(R)\in\mathbb{N}\,s.t.\,\exists\{x^{j}_{R,i}:i=1,\ldots,N(R)\}\,\,s.t.\,\,X_{j}\subset\bigcup_{i=1}^{N(R)}B(x^{j}_{R,i},R)$
and have uniform bounds on their Lipschitz one functions
(3) $\tau_{j}(X_{j})\subset[0,\tau_{max}],$
then a subsequence converges in the intrinsic timed-Hausdorff sense to a compact timed metric space $(X_{\infty},d_{\infty},\tau_{\infty})$. In fact, there exist distance and time preserving timed-Fréchet maps,
(4) $\varphi_{j}\colon X_{j}\to[0,\tau_{max}]\times Z\subset\ell^{\infty},$
where $Z$ is compact, such that
(5) $d_{H}^{[0,\tau_{max}]\times Z}(\varphi_{j}(X_{j}),\varphi_{\infty}(X_{\infty}))\to 0.$
We remark that there are special classes of space-times where the cosmological time functions are easy to control. These include the big-bang space-times and future developed space-times. Sormani–Vega [29] and Sakovich–Sormani [26] defined convergence for such space-times based on pointed GH convergence of Gromov [18] and GH Convergence for Metric Pairs of Che, Galaz-García, Guijarro, and Membrillo Solis [7]. Such notions already have compactness theorems proven by Gromov [18] and by Che and Ahumada Gómez [1] respectively. The intrinsic timed-Hausdorff distance applies to a far wider class of space-times including sub-space-times that exhaust asymptotically flat space-times [26].
It should be noted that some authors have used the Sormani–Vega null distance of [30] to convert space-times into metric spaces without keeping track of the time functions. See the work of Allen, Burtscher, García-Heveling, Kunzinger, and Steinbauer in [2, 3, 6, 20]. Furthermore, Allen, Burtscher, Kunzinger, and Steinbauer [2, 3, 20] have applied Gromov’s original compactness theorem to obtain a GH limit space which is only a metric space with no causal structure unlike our timed metric spaces.
Other notions of weak convergence of space-times do not involve the conversion of the space-times into metric spaces or timed-metric-spaces. Minguzzi–Suhr [22] have a compactness result in the setting of “Lorentzian metric spaces” (which are not metric spaces) building upon work of Noldus [24, 25]. See also the work of Mondino–Sämann [23] in the setting of “Lorentzian pre-length spaces”, as introduced in [19], where they define a notion of Lorentzian Gromov–Hausdorff convergence and obtain a precompactness result using causal diamonds and Lorentzian distances. It is not yet clear how their notions of convergence relate to the Sakovich–Sormani notions defined using timed-metric-spaces in [26]. This is worth exploring further.
We note that, in upcoming work, Sakovich–Sormani will be introducing timed intrinsic flat convergence of space-times whose limit spaces are rectifiable timed-metric-spaces with biLipchitz charts [27]. Our compactness theorem proven in this paper will be applied in that one in an essential way.
The structure of the paper is as follows. In Section 2 we review preliminary notions and results about the Gromov–Hausdorff distance, the classical Gromov’s compactness theorem, and the intrinsic timed-Hausdorff distance. Our compactness theorem is proven in Sections 3.13.3. In Section 3.1 we begin the proof by imitating Gromov’s proof of his compactness theorem in [16]. As we added significant details to his original proof, we developed the notion of “addresses” for points in Section 3.2. With these “addresses” we complete a stronger statement of the compactness theorem which controls the convergence of time in a uniform way in Section 3.3.
Our notion of “addresses” provides a new way of stating Gromov’s Compactness Theorem for sequences of metric spaces, that we state and prove in Section 4.1. We also restate an Arzelà–Ascoli Theorem for sequences of real valued uniformly bounded Lipschitz functions on a converging sequence of metric spaces using “addresses” in Section 4.2. We suggest a more general Arzelà–Ascoli Conjecture that could be proven using “addresses” in Section 4.3.
Acknowledgements: We would like to thank Anna Sakovich (Uppsala) for online conversations in July 2025 that lead us to pursuing this project. We all gratefully acknowledge support from the Simons Center for Geometry and Physics and Stony Brook University where we had the opportunity to meet together in person in September 2025 as part of the Program Geometry and Convergence in Mathematical General Relativity. R. Perales was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/EFP6]. M. Che was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/STA32]. C. Sormani was funded in part by a PSC-CUNY grant. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
2. Background
2.1. Review of the Hausdorff Distance
The Hausdorff distance between subsets, $U,W\subset Z$, lying in a common metric space, $(Z,d_{Z})$, is defined by
(6) $d_{H}^{Z}\Big(U,W\Big)=\inf\left\{r>0:\begin{array}[]{c}\forall u\in U,\exists w\in W\textrm{ s.t. }d_{Z}(u,w)<r\\ \forall w\in W,\exists u\in U\textrm{ s.t. }d_{Z}(u,w)<r\end{array}\right\}.$
Recall that the Hausdorff distance between sets, $U,W\subset Z$, is extrinsic in the sense that it depends on the distance in the extrinsic space.
2.2. Review of Gromov–Hausdorff Distance
Gromov defined his intrinsic Hausdorff distance between metric spaces, $(X_{j},d_{j})$, which do not lie in a common metric space in [18]. This is now called the Gromov–Hausdorff (GH) distance. The GH distance is intrinsic in the sense that it depends only on the intrinsic geometry of the two given metric spaces and not the extrinsic geometry of some common ambient space.
In order to define this distance intrinsically, Gromov considered all possible common metric spaces, $(Z,d_{Z})$, and considered maps: $\varphi_{j}\colon X_{j}\to Z$ that are distance preserving
(7) $d_{Z}(\varphi_{j}(p),\varphi_{j}(q))=d_{j}(p,q)\qquad\forall p,q\in X_{j}.$
and takes the infimum:
(8) $d_{GH}\bigg((X_{1},d_{1}),(X_{2},d_{2})\bigg)=\inf d_{H}^{Z}\bigg(\varphi_{1}(X_{1}),\varphi_{2}(X_{2})\bigg).$
Thanks to the infimum taken over all possible distance preserving maps, the Gromov–Hausdorff distance depends only on the intrinsic geometry of the original pair of metric spaces not on any particular external metric space, $(Z,d_{Z})$. Yet at the same time the GH distance between metric spaces has many of the same properties as the Hausdorff distance. In particular, there are correspondences between $X_{1}$ and $X_{2}$ with controlled distance distortion that are often used to define the GH distance, but we will not use correspondences in this paper.
2.3. Review of Gromov’s Compactness Theorem
Recall Gromov’s Compactness Theorem proven in [18] and [16]:
Theorem 2.1 (Gromov Compactness Theorem).
If $(X_{j},d_{j})$ is a sequence of compact metric spaces that are equibounded,
(9) $\exists D>0\,s.t.\,\operatorname{diam}_{d_{j}}(X_{j})\leq D,$
equicompact,
(10) $\forall R\in(0,D]\,\,\exists N(R)\in\mathbb{N}\,s.t.\,\exists x^{j}_{R,i}\in X_{j}\,\,s.t.\,\,X_{j}\subset\bigcup_{i=1}^{N(R)}B(x^{j}_{R,i},R),$
then a subsequence converges in the Gromov–Hausdorff sense to a compact metric space $(X_{\infty},d_{\infty})$:
(11) $d_{GH}((X_{j},d_{j}),(X_{\infty},d_{\infty}))\to 0.$
In fact, there exist a compact metric space, $(Z,d_{Z})$, and distance preserving maps,
(12) $\varphi_{j}\colon X_{j}\to Z,$
such that the Hausdorff distance between the images converges to zero:
(13) $d_{H}^{Z}\bigg(\varphi_{j}(X_{j}),\varphi_{\infty}(X_{\infty})\bigg)\to 0.$
2.4. Review of Fréchet Maps
Here we review the notion of a Fréchet map [12].  [^2] In v1 of [26] this was accidentally attributed to Kuratowski. The actual definition of a Kuratowski map can be consulted here [21].
First recall that a separable metric space, $(X,d)$, is a metric space containing a countable dense collection of points.
Recall also the Banach space
[…]
where
(15) $d_{\ell^{\infty}}((s_{1},s_{2},...),(r_{1},r_{2},...))=\sup\{|s_{i}-r_{i}|\,:\,i\in{\mathbb{N}}\}.$
Definition 2.2.
Given a separable metric space, $(X,d)$, for any countable dense collection of points $\mathcal{N}=\{x_{1},x_{2},$...\} of $X$, one can define a Fréchet map
(16) $\kappa_{X}=\kappa_{X,\mathcal{N}}\colon(X,d)\to(\ell^{\infty},d_{\ell^{\infty}})$
by
(17) $\kappa_{X}(x)=(d(x_{1},x),d(x_{2},x),\ldots)\subset\ell^{\infty}.$
Now we state Fréchet’s embedding theorem [13]:
Theorem 2.3 (Fréchet).
Fréchet maps as in Definition 2.2 are distance preserving:
(18) $d_{\ell^{\infty}}(\kappa_{X}(x),\kappa_{X}(y))=d(x,y)\quad\forall\,x,y\in X.$
It is well known that one can define a version of the Gromov–Hausdorff distance by taking the infimum of the Hausdorff distance between images of Fréchet maps into $\ell^{\infty}$.
Definition 2.4.
The Fréchet-GH distance between two metric spaces is defined as
(19) $d_{\kappa-GH}((X,d_{X}),(Y,d_{Y}))=\inf d_{H}^{\ell^{\infty}}(\kappa_{X}(X),\kappa_{Y}(Y))$
where the infimum is over all pairs of Fréchet maps
[…]
that is, the infimum is taken over all selections of countably dense points and reorderings of these selected points in $X$ and in $Y$.
By Fréchet’s embedding theorem, for any pair of compact metric-spaces one immediately has
(21) $d_{\kappa-GH}((X,d_{X}),(Y,d_{Y}))\geq d_{GH}((X,d_{X}),(Y,d_{Y})).$
In fact, this notion is biLipschitz equivalent to the Gromov–Hausdorff distance:
(22) $d_{GH}((X,d_{X}),(Y,d_{Y}))\leq d_{\kappa-GH}((X,d_{X}),(Y,d_{Y}))\leq 2d_{GH}((X,d_{X}),(Y,d_{Y})).$
This close relationship with the Gromov–Hausdorff distance is discussed and reproven in the background of [26] and directly inspired their definition of the intrinsic timed-Hausdorff distance using timed-Fréchet maps.
2.5. Review of Timed-Metric-Spaces
The notion of a timed-metric-space was introduced by Sakovich–Sormani in [26] because they had constructed a canonical way to convert certain classes of smooth Lorentzian space-times into timed-metric-spaces keeping track of the causal structure of the original Lorentzian manifold. For our purposes, we need only the abstract notion of a timed-metric-space:
Definition 2.5.
[…]
(23) $\forall q\in X\quad p\in J_{X}^{+}(q)\Longleftrightarrow\tau(p)-\tau(q)=d(p,q).$
Given a map $F\colon(X,d_{X},\tau_{X})\to(Y,d_{Y},\tau_{Y})$ we say that it is distance preserving iff
(24) $d_{Y}(F(p),F(q))=d_{X}(p,q)\qquad\forall p,q\in X.$
We say that $F$ is time preserving iff
(25) $\tau_{Y}(F(p))=\tau_{X}(p)\qquad\forall p\in X.$
We say that $F$ preserves the causal structure iff
(26) $F(p)\in J^{+}_{Y}(F(q))\Longleftrightarrow p\in J^{+}_{X}(q)$
which happens when $F$ is both distance and time preserving.
In this paper we focus on the geometry of timed-metric-spaces and the intrinsic timed-Hausdorff distance. Our work applies to study Lorentzian manifolds and their intrinsic timed-Hausdorff limits, but we will not be exploring this within the paper. Hence, we will not refer further to smooth manifolds.
2.6. Review of Timed-Fréchet Maps
Sakovich–Sormani introduced the following notion in [26, Definition 4.21]:
Definition 2.6.
Given a compact timed-metric-space, $(X,d,\tau)$, and a countably dense collection of points,
(27) ${\mathcal{N}}=\{x_{1},x_{2},x_{3},\ldots\}\subset X,$
we define a timed-Fréchet map [^4] In v1 of [26] these were accidentally named timed-Fréchet maps and this will be fixed. $\kappa_{\tau,X}=\kappa_{\tau,X,{\mathcal{N}}}\colon X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty}$, by
[…]
For the reader’s convenience, we state here propositions 4.24 and 4.25 from [26].
Proposition 2.7.
Any timed-Fréchet map, $\kappa_{\tau,X}\colon X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty},$ is distance preserving,
(29) $d_{\ell^{\infty}}(\kappa_{\tau,X}(x),\kappa_{\tau,X}(y))=d(x,y),$
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau\colon X\to[0,\tau_{max}]$ and any choice of countably dense points in $X$.
Proposition 2.8.
If we consider the infinite dimensional timed-metric-space
(30) $([0,\tau_{max}]\times\ell^{\infty},d_{\ell^{\infty}},\tau_{\infty}),$
where
(31) $\tau_{\infty}(w_{0},w_{1},w_{2},\ldots)=w_{0},$
and endowed with a causal structure defined by
(32) $\forall z\in[0,\tau_{max}]\times\ell^{\infty}\quad w\in J_{\infty}^{+}(z)\Longleftrightarrow\tau_{\infty}(w)-\tau_{\infty}(z)=d_{\ell^{\infty}}(w,z),$
then every timed-Fréchet map $\kappa_{\tau,X}\colon X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty}$ is time preserving:
(33) $\tau_{\infty}(\kappa_{\tau,X}(x))=\tau(x)$
and preserves causality:
(34) $\kappa_{\tau,X}(p)\in J_{\infty}^{+}(\kappa_{\tau,X}(q))\Longleftrightarrow p\in J_{X}^{+}(q)$
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau\colon X\to[0,\tau_{max}]$ that encodes causality as in (23).
2.7. Review of the Intrinsic Timed-Hausdorff Distance
In [26, Definition 4.22] Sakovich and Sormani introduced the following notion within their definition of the intrinsic timed-Hausdorff distance between causally null compactifiable space-times:
Definition 2.9.
The intrinsic timed-Hausdorff distance between two compact timed-metric-spaces, $(X_{i},d_{i},\tau_{i})$, $i=1,2$, is given by
(35) $d^{\kappa}_{\tau-H}\Big((X_{1},d_{1},\tau_{1}),(X_{2},d_{2},\tau_{2})\Big)=\inf\,d^{\ell^{\infty}}_{H}(\kappa_{\tau_{1},X_{1}}(X_{1}),\kappa_{\tau_{2},X_{2}}(X_{2}))$
where the infimum is taken over all possible timed-Fréchet maps, $\kappa_{\tau_{1},X_{1}}\colon X_{1}\to\ell^{\infty}$ and $\kappa_{\tau_{2},X_{2}}\colon X_{2}\to\ell^{\infty}$ which are found by considering all countable dense collections of points in the $X_{i}$ and all reorderings of these collections of points.
They proved this notion of distance is definite in [26, Theorem 5.14], which we state below.
Theorem 2.10.
The intrinsic timed-Hausdorff distance between two compact timed-metric-spaces, as in definition 2.9, is definite in the sense that:
(36) $d^{\kappa}_{\tau-H}\Big((X,d_{X},\tau_{X}),(Y,d_{Y},\tau_{Y})\Big)=0$
if and only if there is a distance and time preserving bijection
(37) $F\colon(X,d_{X},\tau_{X})\to(Y,d_{Y},\tau_{Y}).$
3. Proving the Compactness Theorem with Addresses
In this section we prove our compactness theorem:
Theorem 3.1 (Timed Gromov Compactness Theorem).
If $(X_{j},d_{j},\tau_{j})$ is a sequence of compact timed metric spaces that are equibounded,
(38) $\exists D>0\,s.t.\,\operatorname{diam}_{d_{j}}(X_{j})\leq D,$
equicompact,
(39) $\forall R\in(0,D]\,\exists N(R)\in\mathbb{N}\,s.t.\,\exists\{x^{j}_{R,i}:i=1,\ldots,N(R)\}\,\,s.t.\,\,X_{j}\subset\bigcup_{i=1}^{N(R)}B(x^{j}_{R,i},R)$
and have uniform bounds on their Lipschitz one functions
(40) $\tau_{j}(X_{j})\subset[0,\tau_{max}],$
then a subsequence converges in the intrinsic timed-Hausdorff sense to a compact timed metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. In fact, up to a subsequence, there exist distance and time preserving timed-Fréchet maps,
(41) $\varphi_{j}\colon X_{j}\to[0,\tau_{max}]\times Z\subset\ell^{\infty},$
where $Z$ is compact, such that
(42) $d_{H}^{[0,\tau_{max}]\times Z}(\varphi_{j}(X_{j}),\varphi_{\infty}(X_{\infty}))\to 0.$
In addition, there is an uncountable index set $\mathcal{A}$, and surjective index maps,
(43) ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}\quad\textrm{ and }\quad{\mathcal{I}}^{\infty}\colon{\mathcal{A}}\to\bar{X}_{\infty}$
such that we have uniform convergence as $j\to\infty$,
(44) $\sup_{\alpha\in{\mathcal{A}}}d_{\ell^{\infty}}\Big(\varphi_{j}({\mathcal{I}}^{j}(\alpha)),\varphi_{\infty}({\mathcal{I}}^{\infty}(\alpha))\Big)\to 0.$
In particular, we have uniform convergence of time,
(45) $\sup_{\alpha\in{\mathcal{A}}}|\tau_{j}({\mathcal{I}}^{j}(\alpha))-\tau_{\infty}({\mathcal{I}}^{\infty}(\alpha))|\to 0,$
and uniform convergence of distances,
(46) $\sup_{\alpha,\alpha^{\prime}\in{\mathcal{A}}}|d_{j}({\mathcal{I}}^{j}(\alpha),{\mathcal{I}}^{j}(\alpha^{\prime}))-d_{\infty}({\mathcal{I}}^{\infty}(\alpha),{\mathcal{I}}^{\infty}(\alpha^{\prime}))|\to 0.$
3.1. Selection of Balls in a Compact Metric Space
The following proposition is from the first part of Gromov’s proof of his Compactness Criterion in [16]. Since we will apply it later to prove our Theorem 3.1 we provide more details in the proof and add in the precise indexing of the points $x_{a}$. The proposition allows us to carefully select a countable dense collection of points covering each metric space and indexing them in a way which keeps track of where each point is located. See Figure 3.1. We will apply this indexing later to define the index set of “addresses” $\mathcal{A}$.
Proposition 3.2.
Suppose $X_{j}$ is a metric space satisfying the equiboundedness and equicompactness conditions as in the statement of Theorem 2.1. Take the sequence $\varepsilon_{i}=2^{-i}$ and let $N_{i}=N(\varepsilon_{i})$ be natural numbers such that $X_{j}$ can be covered by $N_{i}$ balls of radius $\varepsilon_{i}$. Then we can find a countable dense subset
(47) $\{x_{a}^{j}:\,a\in A\}\subset X_{j}$
using a uniform index set $A$, not depending on $j$, as described below:
For each $i\in{\mathbb{N}}$, we let $A_{i}$ be the finite set,
(48) $A_{i}=\Bigl\{(a_{1},a_{2},\ldots,a_{i}):\,1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2},\dots,1\leq a_{i}\leq N_{i}\Bigr\},$
and we define the projection maps,
(49) $p_{i}\colon A_{i+1}\to A_{i}\qquad p_{i}\colon(a_{1},\ldots,a_{i+1})\mapsto(a_{1},\ldots,a_{i}).$
Then we define the countable index set for our countable collection of points as
(50) $A=\bigsqcup_{i=1}^{\infty}A_{i}.$
We will define maps
(51) $I^{j}_{i}\colon A_{i}\to X_{j}\quad\textrm{ and }\quad I^{j}\colon A\to X_{j},$
where $I^{j}$ is given by
(52) $I^{j}(a)=I^{j}_{i}(a)\in X_{j}\qquad\forall a\in A_{i}\subset A,$
and set
(53) $x^{j}_{a}=I^{j}(a)=I^{j}_{i}(a)\qquad\forall a\in A_{i},$
which will be well chosen centers of balls of radius $\varepsilon_{i}$ in $X_{j}$ for $a\in A_{i}$. Specifically, first $I^{j}_{i}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, which means
(54) $X_{j}\subset\bigcup_{a\in A_{i}}B_{d_{j}}(x^{j}_{a},\varepsilon_{i})$
Second, for each $a\in A_{i+1}$,
(55) $I^{j}_{i+1}(a)\in B_{d_{j}}(I^{j}_{i}(p_{i}(a)),2\varepsilon_{i})\subset X_{j}$
which means
(56) $x^{j}_{(a_{1},...a_{i},a_{i+1})}\in B_{d_{j}}(x^{j}_{(a_{1},...a_{i})},2\varepsilon_{i})$
for any $a=(a_{1},\dots,a_{i+1})\in A_{i+1}$.
Hence, the collection of points
(57) ${\mathcal{N}}_{j}=I^{j}(A)=\{x_{a}^{j}:\,a\in A\}=\bigcup_{i\in{\mathbb{N}}}I^{j}_{i}(A_{i})\subset X_{j}$
is countably dense in $X_{j}$ and has been carefully indexed to satisfy the previous two properties.
Figure 3.1. An index set with $N(1)=1,N(2)=3,N(3)=4$ and the corresponding points in $X_{j}$.
Proof.
We will define $A_{i}$, $p_{i}$ and $I^{j}_{i}$ satisfying (54) and (56) by induction on $i\in{\mathbb{N}}$. Note that, once this induction is completed, we can see that the set $I^{j}(A)$ is dense in $X_{j}$ because we can take any $x\in X_{j}$ and any $\delta>0$, and choose $i$ sufficiently large that $\varepsilon_{i}<\delta$, and then apply (54) to see that,
(58) $\exists x^{j}_{a}\in I^{j}(A)\,\,s.t.\,\,d_{j}(x,x^{j}_{a})<\varepsilon_{i}<\delta.$
By construction, the set $A$ and its image $I^{j}(A)$ are countable because they are countable unions of finite sets, $A_{i}$, and their images, $I^{j}_{i}(A_{i})$, respectively. Thus we need only complete the induction to prove this proposition.
We begin the induction argument with $i=1$. By equiboundedness and the definition of $N_{1}=N(\varepsilon_{1})$, we have an $\varepsilon_{1}$-net of points,
(59) $\{x^{j}_{(1)},x^{j}_{(2)},x^{j}_{(3)},...x^{j}_{(N_{1})}\}\subset X$
indexed by
(60) $A_{1}=\{(1),(2),\ldots,(N_{1})\}$
such that
(61) $X_{j}\subset\bigcup_{a\in A_{1}}B_{d_{j}}(x^{j}_{a},\varepsilon_{1})=\bigcup_{a_{1}=1}^{N_{1}}B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})$
which gives us (54) for $i=1$.
By equiboundedness and the definition of $N_{2}=N(\varepsilon_{2})$, for each ball of radius $\varepsilon_{1}$ (indexed by $a_{1}=1,2,...,N_{1}$) as above, we can choose $N_{2}$ points in $X_{j}$ such that
(62) $B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})\subset\bigcup_{a_{2}=1}^{N_{2}}B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2})$
with
(63) $B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})\cap B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2})\neq\emptyset\qquad a_{2}=1,\ldots,N_{2}.$
Thus
(64) $x^{j}_{(a_{1},a_{2})}\in B_{d_{j}}(x^{j}_{(a_{1})},2\varepsilon_{1})\subset X_{j},$
So we have an index set,
(65) $A_{2}=\{(a_{1},a_{2}):1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2}\},$
and a projection map
(66) $p_{1}\colon A_{2}\to A_{1}\qquad p_{1}\colon(a_{1},a_{2})\mapsto(a_{1})$
and the map $I^{j}_{2}\colon A_{2}\to X_{j}$ identifying the center points,
(67) $x^{j}_{(a_{1},a_{2})}=I^{j}_{2}((a_{1},a_{2}))$
of the balls of radius $\varepsilon_{2}$. Observe that by (64) and the definition of the projection map, $p_{1}$, we have indexed the points
(68) $x^{j}_{(a_{1},a_{2})}\in B_{d_{j}}(x^{j}_{p_{1}(a_{1},a_{2})},2\varepsilon_{1})\subset X_{j},$
which gives (56) for $i=1$. This completes the base step of our proof by induction.
For further understanding, note that by combining (61) and (62) we have
$X_{j}\subset\bigcup_{a_{1}=1}^{N_{1}}B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})\subset\bigcup_{a_{1}=1}^{N_{1}}\bigcup_{a_{2}=1}^{N_{2}}B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2})=\bigcup_{a\in A_{2}}B_{d_{j}}(x^{j}_{a},\varepsilon_{2})$
which gives us (54) for $i=2$.
Now we proceed with the induction step. Assume we already have a collection of points indexed by the sets
$A_{1}$ $=$ $\{(a_{1}):1\leq a_{1}\leq N_{1}\},$
$A_{2}$ $=$ $\{(a_{1},a_{2}):1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2}\},$
$A_{3}$ $=$ $\{(a_{1},a_{2},a_{3}):1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2},\,1\leq a_{3}\leq N_{3}\},$
$\vdots$
$A_{i}$ $=$ $\{(a_{1},a_{2},\ldots,a_{i}):\,1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2},\ldots,1\leq a_{i}\leq N_{i}\},$
with projection maps,
(69) $p_{1}\colon A_{2}\to A_{1},\quad p_{2}\colon A_{3}\to A_{2},\quad\ldots,\quad p_{i-1}\colon A_{i}\to A_{i-1},$
and maps
(70) $I^{j}_{1}\colon A_{1}\to X_{j},\quad I^{j}_{2}\colon A_{2}\to X_{j},\quad\ldots,\quad I^{j}_{i}\colon A_{i}\to X_{j},$
identifying the center points,
$x^{j}_{(a_{1},a_{2})}\,\,=$ $I^{j}_{2}((a_{1},a_{2}))$ $\in\,\,B_{d_{j}}(x^{j}_{(a_{1})},2\varepsilon_{1}),$
$x^{j}_{(a_{1},a_{2},a_{3})}\,\,=$ $I^{j}_{3}((a_{1},a_{2},a_{3}))$ $\in\,\,B_{d_{j}}(x^{j}_{(a_{1},a_{2})},2\varepsilon_{2}),$
$\vdots$
$x^{j}_{(a_{1},a_{2},...,a_{i})}\,\,=$ $I^{j}_{i}((a_{1},a_{2},...,a_{i}))$ $\in\,\,B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i-1})},2\varepsilon_{i-1}),$
such that
$B_{d_{j}}(x^{j}_{(a_{1})},\varepsilon_{1})$ $\subset$ $\bigcup_{a_{2}=1}^{N_{2}}B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2}),$
$B_{d_{j}}(x^{j}_{(a_{1},a_{2})},\varepsilon_{2})$ $\subset$ $\bigcup_{a_{3}=1}^{N_{3}}B_{d_{j}}(x^{j}_{(a_{1},a_{2},a_{3})},\varepsilon_{3}),$
$\vdots$
$B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i-1})},\varepsilon_{i-1})$ $\subset$ $\bigcup_{a_{i}=1}^{N_{i}}B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i-1},a_{i})},\varepsilon_{i}).$
By equicontinuity, each ball of radius $\varepsilon_{i}$ centered at
(71) $x^{j}_{(a_{1},a_{2},...,a_{i})}=I^{j}_{i}((a_{1},a_{2},...,a_{i})),$
can be covered by $N_{i+1}$ balls of radius $\varepsilon_{i+1}$ centered at points
(72) $x^{j}_{(a_{1},a_{2},...,a_{i},a_{i+1})}\,\,\in\,\,B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i-1},a_{i})},2\varepsilon_{i}),$
where $a_{i+1}\in\{1,2,...,N_{i+1}\}$. That is
(73) $B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i})},\varepsilon_{i})\,\,\subset\,\,\bigcup_{a_{i+1}=1}^{N_{i+1}}B_{d_{j}}(x^{j}_{(a_{1},a_{2},...,a_{i},a_{i+1})},\varepsilon_{i+1}).$
So we have all these new points together indexed by the new index set,
(74) $A_{i+1}=\{(a_{1},a_{2},\ldots,a_{i+1}):1\leq a_{1}\leq N_{1},\,1\leq a_{2}\leq N_{2},\ldots,1\leq a_{i+1}\leq N_{i+1}\},$
and new maps $I^{j}_{i+1}\colon A_{i+1}\to X_{j}$ defined by
(75) $x^{j}_{(a_{1},a_{2},...,a_{i},a_{i+1})}=I^{j}_{i+1}((a_{1},a_{2},...,a_{i},a_{i+1})).$
The new projection map,
(76) $p_{i}\colon A_{i+1}\to A_{i},$
with (72) gives (56) and with (73) gives (54) inductively. ∎
3.2. The Index Set of Addresses
A natural consequence of our description of Gromov’s Proposition 3.2, is that we have a uniform way of providing an “addresses” to all points in equicompact collections of metric spaces. More precisely, we have the following definition and proposition.
Definition 3.3.
Given a set $A$ and maps $p_{i}$ as in Proposition 3.2, an address is a sequence
(77) $\alpha=\{\alpha_{1},\alpha_{2},\alpha_{3},...\}\subset A$
such that
(78) $\alpha_{i}=(a_{1},a_{2},...a_{i})\in A_{i}\subset A\qquad\forall i\in{\mathbb{N}},$
satisfying
(79) $p_{i}(\alpha_{i+1})=\alpha_{i}\qquad\forall i\in{\mathbb{N}}.$
We denote by $\mathcal{A}$ the set of all such addresses (see Figure 3.2).
Proposition 3.4.
Let
(80) $q_{i}\colon{\mathcal{A}}\to A_{i}\textrm{ s.t. }q_{i}(\alpha)=\alpha_{i}.$
Then for any $X_{j}$ and any maps $I^{j}_{i}\colon A_{i}\to X_{j}$ satisfying (51)- (55) of Proposition 3.2, the map, ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}$, defined by
(81) ${\mathcal{I}}^{j}(\alpha)=\lim_{i\to\infty}I^{j}_{i}(q_{i}(\alpha))$
is well defined and surjective. That is, every point $x\in X_{j}$, has an address,
(82) $\alpha=\alpha^{j}(x)=\{\alpha^{j}_{1}(x),\alpha^{j}_{2}(x),...\}\in{\mathcal{A}},$
such that $x={\mathcal{I}}^{j}(\alpha^{j}(x))$. In fact,
(83) $x\in\bar{B}_{d_{j}}(I^{j}_{i}(\alpha^{j}_{i}(x)),\varepsilon_{i})\qquad\forall i\in{\mathbb{N}}.$
We do not claim this address is unique. Moreover, we get the same result even if (54) and (55) hold for closed balls rather than open balls.
Figure 3.2. An address $\alpha$ identifying a sequence of points in $X_{j}$ approaching the point ${\mathcal{I}}^{j}(\alpha)$.
Proof.
To prove ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}$ is well defined, we need only show the sequence of points,
(84) $\{I^{j}_{i}(\alpha_{i})\colon\,i\in\mathbb{N}\}$
is Cauchy. By $p_{i}(\alpha_{i+1})=\alpha_{i}$, and the property (55) for $I^{j}_{i}$ proven in Proposition 3.2, we have
(85) $I^{j}_{i+1}(\alpha_{i+1})\in\bar{B}_{d_{j}}(I^{j}_{i}(p_{i}(\alpha_{i+1})),2\varepsilon_{i})=\bar{B}_{d_{j}}(I^{j}_{i}(\alpha_{i}),2\varepsilon_{i}).$
Thus for all $i\in{\mathbb{N}}$, we have
(86) $d_{j}(I^{j}_{i}(\alpha_{i}),I^{j}_{i+1}(\alpha_{i+1}))\leq 2\varepsilon_{i}=1/2^{i-1}.$
Applying the triangle inequality and then this estimate repeatedly we have
$d_{j}(I^{j}_{i}(\alpha_{i}),I^{j}_{i+k}(\alpha_{i+k}))$ $\leq$ $d_{j}(I^{j}_{i}(\alpha_{i}),I^{j}_{i+1}(\alpha_{i+1}))+\cdots+d_{j}(I^{j}_{i+k-1}(\alpha_{i+k-1}),I^{j}_{i+k}(\alpha_{i+k}))$
$\leq$ $\tfrac{1}{2^{i-1}}+\cdots+\tfrac{1}{2^{i+k-2}}<\tfrac{1}{2^{i}}=\varepsilon_{i}.$
So the sequence is Cauchy. Since $X_{j}$ is compact, it is complete and so the limit in (81) exists, defining ${\mathcal{I}}^{j}(\alpha)$.
Next we show surjectivity. Take any $x\in X_{j}$. We then apply property (54). Since $X_{j}$ is covered by balls of radius $\varepsilon_{1}$ we know there exists $(a^{j}_{1}(x))\in A_{1}$ such that
(87) $x\in\bar{B}_{d_{j}}(I^{j}_{1}(a^{j}_{1}(x)),\varepsilon_{1}).$
This ball is covered by balls of radius $\varepsilon_{2}$, with centers indexed by $(a_{1},a_{2})\in A_{2}$ so there exists $a^{j}_{2}(x)\in\{1,\ldots,N_{2}\}$ such that
(88) $x\in\bar{B}_{d_{j}}(I^{j}_{2}((a^{j}_{1}(x),a^{j}_{2}(x))),\varepsilon_{2}).$
Continuing forward we select each $a^{j}_{i}(x)$ such that
(89) $x\in\bar{B}_{d_{j}}(I^{j}_{i}((a^{j}_{1}(x),a^{j}_{2}(x),...,a^{j}_{i}(x))),\varepsilon_{i}).$
Taking
(90) $\alpha^{j}(x)=\{\alpha^{j}_{1}(x),\alpha^{j}_{2}(x),...\}\in\mathcal{A}$
where
(91) $\alpha^{j}_{i}(x)=(a^{j}_{1}(x),a^{j}_{2}(x),...,a^{j}_{i}(x))$
gives us (83). By the definition of ${\mathcal{I}}^{j}$, and then by (83), we have
(92) ${\mathcal{I}}^{j}(\alpha^{j}(x))=\lim_{i\to\infty}I^{j}_{i}(\alpha^{j}_{i}(x))=x.$
3.3. Compactness Theorem with Addresses
We can now prove our compactness theorem.
Proof of Theorem 3.1.
The proof begins as in Gromov’s Compactness Criterion in [16] using the equicompactness as we have stated in Proposition 3.2 to define a countable index set $A$ defined as the disjoint union of finite sets,
(93) $A=\bigsqcup_{i=1}^{\infty}A_{i},$
and a countable dense collection of points
(94) $I^{j}(A)=\{x^{j}_{a}=I^{j}(a):\,a\in A\}\subset X_{j}$
such that each $I^{j}(A_{i})$ is a finite $\varepsilon_{i}=1/2^{i}$ net for $X_{j}$ and has all the other properties described in Proposition 3.2.
Then we follow the proof in [5] of Gromov’s Compactness Theorem, to construct a limit space. By the equiboundedness of the distance, for any $a,b\in A$, we define
(95) $d^{j}_{a,b}=d_{j}(x^{j}_{a},x^{j}_{b})\in[0,D].$
This definition is symmetric:
(96) $d^{j}_{a,b}=d^{j}_{b,a}\qquad\forall a,b\in A;$
and it satisfies the triangle inequality:
(97) $d^{j}_{a,c}\leq d^{j}_{a,b}+d^{j}_{b,c}\qquad\forall a,b,c\in A.$
By the equiboundedness of the time functions, for any $a\in A$, we define
(98) $\tau^{j}_{a}=\tau_{j}(x^{j}_{a})\in[0,\tau_{max}].$
Since the time functions are Lipschitz one, we have
(99) $|\tau^{j}_{a}-\tau^{j}_{b}|\leq d^{j}_{a,b}\qquad\forall a,b\in A.$
By the countability of the index set $A$, we take subsequences of subsequences and diagonalize so that for the final subsequence (which we still index with $j$), we have for each $a\in A$, convergence of the real numbers:
(100) $\lim_{j\to\infty}\tau^{j}_{a}=\tau^{\infty}_{a}\in[0,\tau_{max}].$
We take further subsequences and diagonalize further, so that for the final subsequence (which we still index with $j$), we have for each $a,b\in A$,
(101) $\lim_{j\to\infty}d^{j}_{a,b}=d^{\infty}_{a,b}\in[0,D].$
Taking the limit of (99) we know
(102) $|\tau^{\infty}_{a}-\tau^{\infty}_{b}|\leq d^{\infty}_{a,b}\qquad\forall a,b\in A.$
We now define a pseudo-metric space $(A,d_{\infty})$ with
(103) $d_{\infty}(a,b)=d^{\infty}_{a,b}\in[0,D]$
which satisfies the triangle inequality and is symmetric but is not necessarily definite. By (102) we know that
(104) $d^{\infty}(a,b)=0\,\,\Longrightarrow\,\,\tau^{\infty}_{a}=\tau^{\infty}_{b}.$
Thus we can mod out by points of zero distance apart to define a timed metric space which is definite:
(105) $(X_{\infty},d_{\infty},\tau_{\infty})=(\{[a]:a\in A\},d_{\infty},\tau_{\infty})$
where the points are equivalence classes,
(106) $a]=\{b\in A:\,d_{\infty}(a,b)=0\},$
the distance function is
(107) $d_{\infty}([a],[b])=d^{\infty}_{a,b}\in[0,D],$
and the time function is
(108) $\tau_{\infty}([a])=\tau^{\infty}_{a}\in[0,\tau_{\max}],$
which is well defined by (104) and is Lipschitz one by (102).
We next take the metric completion
(109) $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$
which has the usual definite distance, $d_{\infty}$, extended from $X_{\infty}$ to $\bar{X}_{\infty}$ satisfying symmetry and the triangle inequality, with
(110) $\operatorname{diam}_{d_{\infty}}(\bar{X}_{\infty})\leq D,$
and a bounded time function, $\tau_{\infty}$, extended as a Lipschitz one function from $X_{\infty}$ to $\bar{X}_{\infty}$.
We claim that $\bar{X}_{\infty}$ is compact. Note that to prove this claim is enough to show that for each $i\in\mathbb{N}$,
(111) $\Bigl\{[a]\,:\,a\in A_{i}\Bigr\}\qquad\text{is an $\varepsilon_{i}$-net of $X_{\infty}$.}$
Take $b]\in X_{\infty}$. Thus, there exists $k\in\mathbb{N}$ such that $b\in A_{k}$. Since for each $j\in\mathbb{N}$ we know that $I_{i}^{j}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, by the pigeon-hole principle, there exists $a\in A_{i}$ such that
(112) $d_{j}(I_{i}^{j}(a),I_{i}^{k}(b))\leq\varepsilon_{i}$
for infinitely many indices $i$. Thus, up to a subsequence,
(113) $d_{\infty}([b],[a])=\lim_{j\to\infty}d_{j}(x_{a}^{j},x_{b}^{j})\leq\varepsilon_{i}.$
Hence, we have (111).
Up to here we have followed the proof from [5] of Gromov’s Compactness Theorem, with corresponding additions to address the fact that we have time functions, to construct a compact timed-metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. Since we will show timed-Hausdorff convergence,
(114) $(X_{j},d_{j},\tau_{j})\stackrel{{\scriptstyle\tau-\textrm{H}}}{{\longrightarrow}}(\bar{X}_{\infty},d_{\infty},\tau_{\infty}),$
using timed-Fréchet maps as in Definition 2.9, from now on, we proceed in a different manner to [5].
Note that we can define timed-Fréchet maps,
(115) $\kappa_{\tau_{j},{\mathcal{N}}_{j}}\colon X_{j}\to[0,\tau_{max}]\times\prod_{a\in A}[0,D]\subset\ell^{\infty}$
by
(116) $\kappa_{\tau_{j},{\mathcal{N}}_{j}}(x)=(\tau_{j}(x),\kappa_{{\mathcal{N}}_{j}}(x))$
where
(117) $\kappa_{{\mathcal{N}}_{j}}(x)\in\prod_{a\in A}[0,D]$
has components
(118) $(\kappa_{\mathcal{N}_{j}}(x))_{a}=d_{j}(I^{j}(a),x).$
To define a timed-Fréchet map for $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$, recall that we have a countable index set, $A$, which is the countable disjoint union of finite index sets, $A_{i}$. We define
(119) $I^{\infty}_{i}\colon A_{i}\to X_{\infty}\subset\bar{X}_{\infty}\quad\textrm{ and }\quad I^{\infty}\colon A\to X_{\infty}\subset\bar{X}_{\infty},$
by
(120) $I^{\infty}(a)=I^{\infty}_{i}(a)=[a]\in X_{\infty}\subset\bar{X}_{\infty}\qquad\forall\,a\in A_{i}\subset A,$
and we denote these points as
(121) $x^{\infty}_{a}=I^{\infty}(a)=I^{\infty}_{i}(a).$
Thus,
(122) ${\mathcal{N}}_{\infty}=I^{\infty}(A)=\{x_{a}^{\infty}:\,a\in A\}$
is a countable ordered collection of points, that is dense in $\bar{X}_{\infty}$, since as a set it equals $X_{\infty}$. From this, and recalling (107) and (108), we also have a timed-Fréchet map $\kappa_{\tau_{\infty},\mathcal{N_{\infty}}}$ satisfying (115)-(118) for $j=\infty$.
Before proving (42) we first show that the uniform estimate (44) holds. As a preliminary step we note that we have uniform convergence on finite unions of the $A_{i}$: Recall that by (95), (101) and (107), for each $a,b\in A$, we have pointwise convergence,
(123) $d_{\infty}(x^{\infty}_{a},x^{\infty}_{b})=\lim_{j\to\infty}d^{j}(x^{j}_{a},x^{j}_{b})\in[0,D].$
By (98), (100) and (108) for each $a\in A$, we have pointwise convergence,
(124) $\tau_{\infty}(x^{\infty}_{a})=\lim_{j\to\infty}\tau_{j}(x^{j}_{a}).$
Now, since each $A_{i}$ is finite, we do obtain uniform convergence on finite unions of the $A_{i}$:
(125) $\forall\epsilon>0,\exists N_{\epsilon}^{i}\in{\mathbb{N}},$
such that for all $j\geq N_{\epsilon}^{i}$ we have
(126) $\max\{|\tau_{\infty}(x^{\infty}_{a})-\tau_{j}(x^{j}_{a})|,|d_{\infty}(x^{\infty}_{a},x^{\infty}_{b})-d^{j}(x^{j}_{a},x^{j}_{b})|\}<\epsilon$
where the maximum is over
(127) $a,b\in A_{1}\cup...\cup A_{i}.$
To get better control than just the pointwise convergence as in (123) and (124), we will use Proposition 3.4 to provide an address, $\alpha$, for every point in each $X_{j}$ and $\bar{X}_{\infty}$.
Note that we have proven that $\bar{X}_{\infty}$ satisfies the conclusions (51)-(57), of Proposition 3.2. It satisfies 54 by (111). Moreover, it easily satisfies (55) as follows. Let $a\in A_{i+1}$,
(128) $I^{\infty}_{i+1}(a)\in B_{d_{\infty}}(I^{\infty}_{i}(p_{i}(a)),2\varepsilon_{i})\subset\bar{X}_{\infty}$
We know that for each $j\in\mathbb{N}$,
(129) $d_{j}(x^{j}_{a},x^{j}_{p_{i}(a)})\leq 2\varepsilon_{i}.$
Hence, in the limit, we have
(130) $d_{\infty}(x^{\infty}_{a},x^{\infty}_{p_{i}(a)})=\lim_{j\to\infty}d_{j}(x^{j}_{a},x^{j}_{p_{i}(a)})\leq 2\varepsilon_{i},$
which is the desired estimate.
Since we have applied Proposition 3.2 for each $X_{j}$, $j\in\mathbb{N}$, and the previous paragraph, for each $j\in\mathbb{N}\sqcup\{\infty\}$ there exist surjective maps, ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}$, $j\in\mathbb{N}\cup\{\infty\}$ defined by
(131) ${\mathcal{I}}^{j}(\alpha)=\lim_{i\to\infty}I^{j}_{i}(q_{i}(\alpha))$
that satisfy the conditions of Proposition 3.4.
We claim that for any $\varepsilon>0$ there exist $N_{\varepsilon}\in\mathbb{N}$ such that for all $j\geq N_{\varepsilon}$ we have for any address $\alpha$,
(132) $|\tau_{j}({\mathcal{I}}^{j}(\alpha))-\tau_{\infty}({\mathcal{I}}^{\infty}(\alpha))|\leq\varepsilon$
and that additionally for any $a\in A$, we also have
(133) $|\kappa_{j,a}({\mathcal{I}}^{j}(\alpha))-\kappa_{\infty,a}({\mathcal{I}}^{\infty}(\alpha))|\leq\varepsilon.$
Let $\epsilon=\varepsilon/3$ and $i\in\mathbb{N}$ such that $\varepsilon_{i}<\epsilon$. Then there exists $N_{\epsilon}^{i}\in{\mathbb{N}}$, as in (125), such that for all $j\geq N_{\epsilon}^{i}$ we have (126).
We obtain the first inequality using the triangle inequality,
(134) $|\tau_{j}({\mathcal{I}}^{j}(\alpha))-\tau_{\infty}({\mathcal{I}}^{\infty}(\alpha))|\,\,\leq\,\,$ $|\tau_{j}({\mathcal{I}}^{j}(\alpha))-\tau_{j}(I^{j}_{i}(\alpha_{i}))|$
(135) $+|\tau_{j}(I^{j}_{i}(\alpha_{i}))-\tau_{\infty}(I^{j}_{i}(\alpha_{i}))|$
(136) $+|\tau_{\infty}({\mathcal{I}}^{\infty}(\alpha))-\tau_{\infty}(I^{\infty}_{i}(\alpha_{i}))|$
(137) $\,\,\leq\,\,$ $\,\,d_{j}({\mathcal{I}}^{j}(\alpha),I^{j}_{i}(\alpha_{i}))$
(138) $+|\tau_{j}(I^{j}_{i}(\alpha_{i}))-\tau_{\infty}(I^{j}_{i}(\alpha_{i}))|$
(139) $+d_{\infty}({\mathcal{I}}^{\infty}(\alpha),I^{\infty}_{i}(\alpha_{i}))$
(140) $\leq\,\,$ $\,\,\varepsilon_{i}+\epsilon+\varepsilon_{i}\,\,\leq\,\,\varepsilon\qquad\forall j\geq N^{i}_{\epsilon},\quad\quad\quad$
where for the second and third term in the right hand side we used that $\tau_{j},\tau_{\infty}$ are Lipschitz one and (83), and for the term in the middle we use (126). Similarly, we get the second inequality,
(141) $|\kappa_{j,a}({\mathcal{I}}^{j}(\alpha))-\kappa_{\infty,a}({\mathcal{I}}^{\infty}(\alpha))|\,\,\leq\,\,$ $|\kappa_{j,a}({\mathcal{I}}^{j}(\alpha))-\kappa_{j,a}(I^{j}_{i}(\alpha_{i}))|$
(142) $+\,\,|\kappa_{j,a}(I^{j}_{i}(\alpha_{i}))-\kappa_{\infty,a}(I^{\infty}_{i}(\alpha_{i}))|$
(143) $+\,\,|\kappa_{\infty,a}({\mathcal{I}}^{\infty}(\alpha))-\kappa_{\infty,a}(I^{\infty}_{i}(\alpha_{i}))|$
(144) $\,\,\leq\,\,$ $d_{j}({\mathcal{I}}^{j}(\alpha),I^{j}_{i}(\alpha_{i}))$
(145) $+\,\,|\kappa_{j,a}(I^{j}_{i}(\alpha_{i}))-\kappa_{\infty,a}(I^{\infty}_{i}(\alpha_{i}))|$
(146) $+\,\,d_{\infty}({\mathcal{I}}^{\infty}(\alpha),I^{\infty}_{i}(\alpha_{i}))$
(147) $\,\,\leq\,\,$ $\,\,\varepsilon_{i}+\epsilon+\varepsilon_{i}\leq\varepsilon\qquad\forall j\geq N^{i}_{\epsilon},\quad\quad\quad$
where we used that $\kappa_{j},\kappa_{\infty}$ are Lipschitz one, (83) and (126).
In particular, we have uniform convergence
(148) $\tau_{j}\circ{\mathcal{I}}^{j}\to\tau_{\infty}\circ{\mathcal{I}}^{\infty}\textrm{ on }{\mathcal{A}}$
and
(149) $\kappa_{j,a}\circ{\mathcal{I}}^{j}\to\kappa_{\infty,a}\circ{\mathcal{I}}^{\infty}\textrm{ on }{\mathcal{A}}.$
Hence, (45) and (44) hold, where $\varphi_{j}=(\tau_{j},\kappa_{j})$ and $\varphi_{\infty}=(\tau_{\infty},\kappa_{\infty})$.
Now the correspondences
(150) $\mathcal{C}_{j}=\{(\kappa_{\tau_{j},\mathcal{N}_{j}}({\mathcal{I}}^{j}(\alpha)),\kappa_{\tau_{\infty},\mathcal{N}_{\infty}}({\mathcal{I}}^{\infty}(\alpha))):\alpha\in\mathcal{A}\}$
between $\kappa_{\tau_{j},\mathcal{N}_{j}}(X_{j})$ and $\kappa_{\tau_{\infty},\mathcal{N}_{\infty}}(\bar{X}_{\infty})$, for all $j\in\mathbb{N}$, and the uniform convergence (44) imply that
(151) $d^{\ell^{\infty}}_{H}(\kappa_{\tau_{j},\mathcal{N}_{j}}(X_{j}),\kappa_{\tau_{\infty},\mathcal{N}_{\infty}}(\bar{X}_{\infty}))\leq\|\varphi_{j}({\mathcal{I}}^{j}(\alpha))-\varphi_{\infty}({\mathcal{I}}^{\infty}(\alpha))\|_{\infty}\to 0.$
Finally, let us define
(152) $Z=\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})\cup\bigcup_{j\in\mathbb{N}}\kappa_{\mathcal{N}_{j}}(X_{j})\subset\prod_{a\in A}[0,D]\subset\ell^{\infty},$
so $\kappa_{\tau_{j},\mathcal{N}_{j}}\colon X_{j}\to[0,\tau_{max}]\times Z$ for all $j\in\mathbb{N}\cup\{\infty\}$.
We claim that $Z$ is compact. Let $\{z_{j}\}\subset Z$. If there exists $J\in\mathbb{N}$ such that an infinite number of the $z_{j}$’s is contained in $\kappa_{\mathcal{N}_{J}}(X_{J})$ then by compactness of $\kappa_{\mathcal{N}_{J}}(X_{J})$ there is a further subsequence that converges to a point in $\kappa_{\mathcal{N}_{J}}(X_{J})\subset Z$. Otherwise, assume by passing to a subsequence that $z_{j}\in\kappa_{\mathcal{N}_{j}}(X_{j})$ for all $j\in\mathbb{N}$. By (151), for each $j\in\mathbb{N}$ there exists $w_{j}\in\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})$ such that
(153) $d_{\ell^{\infty}}(z_{j},w_{j})\to 0.$
Since we have shown that $\bar{X}_{\infty}$ is compact, there is a subsequence $\{w_{j_{k}}\}$ that converges to a point $w\in\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})\subset Z$. It follows that $\{z_{j_{k}}\}$ converges to $w$. This completes the proof of $Z$ being compact, and finishes the proof of the theorem. ∎
4. Applications of Addresses
In this final section we restate Gromov’s Compactness Theorem and an Arzelà–Ascoli Theorem in terms of addresses and pose an open question about a more general version of Arzelà-Ascoli.
4.1. Gromov’s Compactness Theorem with Addresses
Theorem 4.1 (Gromov Compactness Theorem with Addresses).
If $(X_{j},d_{j})$ is a sequence of compact metric spaces that are equibounded as in (9) and equicompact as in (10) then a subsequence converges in the Gromov–Hausdorff sense to a compact metric space $(\bar{X}_{\infty},d_{\infty})$. In fact, there exist distance preserving Fréchet maps,
(154) $\kappa_{j}\colon X_{j}\to Z\subset\ell^{\infty},$
where $Z$ is compact, such that
(155) $d_{H}^{Z}(\kappa_{j}(X_{j}),\kappa_{\infty}(X_{\infty}))\to 0.$
In addition, there is an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(156) ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}\quad\textrm{ and }\quad{\mathcal{I}}^{\infty}\colon{\mathcal{A}}\to\bar{X}_{\infty}$
such that
(157) $\sup_{\alpha\in{\mathcal{A}}}d_{\ell^{\infty}}\Big(\kappa_{j}({\mathcal{I}}^{j}(\alpha)),\kappa_{\infty}({\mathcal{I}}^{\infty}(\alpha))\Big)\to 0$
as $j\to\infty$. In particular,
(158) $\sup_{\alpha,\alpha^{\prime}\in{\mathcal{A}}}|d_{j}({\mathcal{I}}^{j}(\alpha),{\mathcal{I}}^{j}(\alpha^{\prime}))-d_{\infty}({\mathcal{I}}^{\infty}(\alpha),{\mathcal{I}}^{\infty}(\alpha^{\prime}))|\to 0.$
This theorem follows exactly as in the proof of our Theorem 3.1 simply by removing all references to the time function. For completeness of exposition, we prove it as a trivial corollary of our theorem.
Proof.
Take $\tau_{j}$ to be constant zero valued function. Then we satisfy the hypotheses of Theorem 3.1. So the conclusions hold. Finally apply
(159) $\varphi_{j}(x)=(\tau_{j}(x),\kappa_{j}(x))=(0,\kappa_{j}(x)).$
4.2. Arzelà–Ascoli Theorem with Addresses
Theorem 4.2 (Arzelà–Ascoli-with-Addresses).
Given a sequence of compact metric spaces:
(160) $(X_{j},d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(X_{\infty},d_{\infty}),$
and uniformly bounded $K$-Lipschitz functions for some $K>0$,
(161) $F_{j}\colon X_{j}\to[0,F_{max}],$
then a subsequence, also denoted $F_{j}$, converges to a bounded $K$-Lipschitz function
(162) $F_{\infty}\colon X_{\infty}\to[0,F_{max}]$
in the following sense: There exist distance preserving Fréchet maps,
(163) $\kappa_{j}\colon X_{j}\to Z\subset\ell^{\infty},$
where $Z$ is compact, such that
(164) $d_{H}^{Z}(\kappa_{j}(X_{j}),\kappa_{\infty}(X_{\infty}))\to 0.$
In addition there exist an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(165) ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}\quad\textrm{ and }\quad{\mathcal{I}}^{\infty}\colon{\mathcal{A}}\to\bar{X}_{\infty}$
such that we have uniform convergence as $j\to\infty$, of the images of the address set under the Fréchet maps,
(166) $\sup_{\alpha\in{\mathcal{A}}}d_{\ell^{\infty}}\Big(\kappa_{j}({\mathcal{I}}^{j}(\alpha)),\kappa_{\infty}({\mathcal{I}}^{\infty}(\alpha))\Big)\to 0,$
and under the uniformly bounded Lipschitz functions
(167) $\sup_{\alpha\in{\mathcal{A}}}|F_{j}({\mathcal{I}}^{j}(\alpha))-F_{\infty}({\mathcal{I}}^{\infty}(\alpha))|\to 0.$
Proof.
Gromov proved that with a compact GH limit, we have an equibounded and equicompact sequence of metric spaces [17]. Setting
(168) $\tau_{j}=\tfrac{1}{K}F_{j}$
we have timed metric spaces, $(X_{j},d_{j},\tau_{j})$ satisfying the hypotheses of Theorem 3.1. The conclusions immediately follow from that theorem and re-applying (168). ∎
4.3. Arzelà–Ascoli Conjecture with Addresses
It would be interesting to apply the address method to more precisely describe the following stronger Arzelà–Ascoli Theorem:
Conjecture 4.3.
Given a pair of Gromov–Hausdorff converging sequences of compact metric spaces:
(169) $X_{j}\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}X_{\infty}\quad\textrm{ and }\quad Y_{j}\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}Y_{\infty}$
and given uniformly bounded Lipschitz $K$ functions for some $K>0$,
(170) $F_{j}\colon X_{j}\to Y_{j},$
then a subsequence converges to a bounded Lipschitz $K$ function
(171) $F_{\infty}\colon X_{\infty}\to Y_{\infty}.$
A proof of this theorem can be found in various places including [28]. However a new proof using the addresses would be able to describe the convergence of the functions more precisely as a kind of uniform convergence.